### v1

# Refactoring Codebases through Library Design

Abstract
Maintainable and general software allows developers to build robust applications efficiently, yet achieving these qualities often requires refactoring specialized solutions into reusable components. This challenge becomes particularly relevant as code agents become increasingly accurate at solving isolated programming problems. We investigate code agents’ capacity to refactor code in ways supporting growth and reusability. We present both a method and a benchmark for refactoring: Librarian, a sample-and-rerank method for generating reusable libraries, and MiniCode, a benchmark where code agents must minimize and refactor multiple independent solutions into a joint library. Compared to state-of-the-art code agents, Librarian achieves strong results on both compression and correctness on MiniCode, obtaining compression rates 1.6-2x better than coding agents while also improving correctness. We open-source our code and benchmark at https://code-refactor.github.io/.
1 Introduction
Writing code is mainly a matter of rewriting code: debugging, refactoring, optimizing, and other activities within the software engineering lifecycle. But poor rewrites incur technical debt, with such debt costing up to \$2 trillion annually [1]. This problem will likely worsen as language models become increasingly responsible for generating code, because they excel at solving isolated programming problems, but their context length demands a myopic view of the codebase. It is therefore valuable to understand not just the ability of language models to solve programming problems, but also their ability to rewrite and refactor code in ways that support growth and reuse.
Effective code refactoring at scale is a design problem. When refactoring codebases, developers must navigate design decisions around concerns such as generality, re-usability, and maintainability. A classic example illustrates this design challenge: Human programmers often create overly-specialized, redundant solutions to similar problems and would benefit from redesigning specialized solutions into a shared library. This consolidation requires careful design decisions about the right level of abstraction — neither too specific nor too general — and appropriate interfaces that balance flexibility with usability.
Here we focus on refactoring multiple code sources into a reusable software library, and pose the following question: To what extent can code agents address this problem, both within human-written codebases, and also in language model-generated code? To answer that question, we develop a new method and a benchmark. This goes beyond past work [2, 3, 4, 5, 6, 7, 8] in library learning that synthesized subroutines across small programs in i.e. $\lambda$-calculus, instead tackling the more naturalistic problem of redesigning large bodies of code written in contemporary high-level languages, such as Python, producing classes, methods, and helper functions in the style of a human-written library. We propose a simple method, Librarian (Figure 1), which samples possible code rewrites then reranks those samples based on criteria designed to capture what it means to have a good refactoring. To generate potential rewrites, we develop methods for clustering pieces of code together that share common structure, so that a succinct prompt can rewrite them jointly into their refactored form.
To evaluate our method and systematically assess the capability of current systems to perform such design-intensive refactorings, we introduce a new benchmark, MiniCode. Existing benchmarks such as SWE-Bench [9], Commit0 [10], and RefactorBench [11] primarily focus on functional correctness and do not directly evaluate an agent’s ability to perform large-scale code consolidation or library design. MiniCode fills this gap by explicitly challenging code agents to minimize and refactor multiple specialized code sources into a unified library. MiniCode requires agents to design libraries that expose a general interface shared across multiple use cases in two domains: competition coding and repository-level Python applications. This involves optimizing not only for functional correctness but also for broader software engineering objectives such as reusability and maintainability, making the task substantially more open-ended than prior benchmarks.
Our results show that state-of-the-art code agents, based on Claude 3.7 Sonnet and o4-mini, struggle to jointly preserve correctness and improve reusability across both domains of MiniCode. In the competition coding domain, our method Librarian improves refactoring quality by 1.89x while also enhancing correctness. However, on the repository-level refactoring, even the strongest agents fail to produce high-quality refactorings, highlighting a substantial gap between current capabilities and the demands of design-oriented code rewriting. Addressing this challenge remains an open and important direction for future research.
Figure 1: Overview of the problem that we study and the general structure of its solutions. Given a collection of different code sources, where a source is either program or repository, we design a modular and reusable library. To do this we cluster together related programs into tuples (left), sample different rewrites using a language model (right), and select the rewrites which optimize various criteria, such as compression, while validating candidate rewrites using test cases (pass rate).
2 Related work
Repo-level language model coding agents.
Recent work has explored the application of language models to repository-level software engineering tasks. Existing benchmarks include SWE-bench [9], which evaluates models on their ability to resolve real-life GitHub issues, and Commit-0 [10], which challenges AI systems to reconstruct entire code repositories from scratch. Such benchmarks primarily evaluate functional correctness via unit tests, without assessing the quality or maintainability of the resulting codebase. RefactorBench [11] takes a step in this direction by benchmarking the ability to follow specific refactoring instructions. Our work differs in that we expect models to perform a more open-ended task: Redesigning code to be more modular and compact by discovering and drawing out reused abstractions.
Library Learning.
Systems which perform library learning research discover shared abstractions across a large number of small programs, which they use to automatically define new subroutines. Systems such as DreamCoder [4], Trove [12], LiLo [7], and REGAL [3] automatically construct such libraries with the goal of making future program synthesis tasks easier to solve, once the learned library is in hand. Our work is closest to REGAL [3], which clusters related code and refactors using language models. However, existing library learning approaches have primarily been demonstrated in small-scale, constrained domains, limiting their applicability to typical software engineering tasks, such as consolidating multiple repositories into cohesive libraries. By framing library learning within the context of realistic, large-scale code repository development, we expand the relevance of library learning to everyday software engineering practice.
Program optimization.
While our goal is to optimize the quality of libraries, other works focus on improving code speed through correctness-preserving transformations [13, 14, 15]. Both forms of program optimization, compression and speed, require more exploration than only optimizing for correctness, as there does not exist a ground-truth answer. Prior work on program optimization benchmarks study code at the file level. We propose a benchmark that transforms programs at a larger scale, across multiple code repositories.
3 Problem Statement
In this section, we propose a refactoring task: Given multiple code sources that contain problem-specific implementations, the goal is to create a cohesive library that captures shared abstractions. This library must reduce the total code size while supporting all original use cases, potentially opening up new use cases as well by mining and formalizing latent shared abstractions. This is accomplished by searching for refactorings that are both correct and simple. Correctness is straightforward to define as the fraction of unit tests passed, but simplicity is more elusive.
One potential measure of simplicity is counting the total number of tokens in the proposed library and refactored code. However, just minimizing program size has obvious failure modes: code should also be natural, elegant, and extensible, which can be in tension with merely finding the shortest program. [^1] Perl Golf is a game where participants attempt to write the shortest Perl program accomplishing a given task. The resulting code is famously incomprehensible, even by the standards of Perl. To address these concerns, we follow prior work in program synthesis [16, 6, 17, 18] and define simplicity as the minimum description length (MDL), or negative log probability under a reference distribution.
Formally, we are given a collection of code sources $\left\{\rho_{n}\right\}_{n=1}^{N}$, and output both a new library $\mathcal{L}$, as well as rewritten refactorings of the original code sources, $\left\{\rho^{\prime}_{n}\right\}_{n=1}^{N}$. We define the pass rate $\tau(\rho_{n})$ as the fraction of unit tests program $\rho_{n}$ passes. In practice we are concerned both with the case where we are refactoring several code sources ($N>1$) and also the case where there is only a single large code source we are refactoring ($N=1$).
We optimize the following objective, which rewards refactorings that pass at least as many tests as the original program and minimize MDL:
(,$\ell\left(\mathcal{L},\left\{\rho_{n}^{\prime}\right\}\right)=\begin{cases}-\log p_{\text{LM}}(\mathcal{L})+\sum_{n}-\log p_{\text{LM}}(\rho_{n}^{\prime}\mid\mathcal{L})&\forall\rho_{n},\tau(\rho_{n})\leq\tau(\rho_{n}^{\prime})\\ \infty&\text{otherwise}\end{cases}$ (1)
where $p_{\text{LM}}(\rho_{n}^{\prime}|\mathcal{L})$ is the probability of the suffix $\rho_{n}^{\prime}$ given the prefix $\mathcal{L}$ under a language model, effectively concatenating the library and the program into one prompt, but only counting the perplexity of the later program tokens.
4 Librarian: Refactoring Code to Create Libraries
This section details our method to compress collections of code sources into libraries, while migrating the code sources to use these shared building blocks. Figure 1 illustrates our method, Librarian.
Librarian follows a simple sample-and-rerank framework to maximize our refactoring objective described in Section 3. It maintains and grows a library of useful functions as part of this objective.
Concretely, our framework follows:
,$\mathcal{L}^{\star},\left\{\rho^{\star}_{n}\right\}=\operatorname*{arg\,min}_{\mathcal{L},\left\{\rho^{\prime}_{n}\right\}\in\textsc{Sample}\left(\left\{\rho_{n}\right\}\right)}\ell\left(\mathcal{L},\left\{\rho^{\prime}_{n}\right\}\right).$ (2)
4.1 Sample with clustering
Meaningful abstractions exist primarily among programs that share some functionality or underlying structure. We perform clustering on the input programs to make groups of programs that likely share structures that can be abstracted into general library functions. Most modern language models cannot be prompted with the entire collection of input programs– even long context models cannot process the entirety of e.g the Linux kernel, and even if they could, it is not clear that such a strategy is the most efficient way of focusing the language model’s attention.
We consider clustering algorithms for discovering small groups of related code; we call these tuples. This extends REGAL [3], which clusters programs solving similar problems by assuming each program is paired with a natural language description of the problem it solves, and clustering embeddings of those descriptions. But programs solving similar problems do not necessarily have similar structure. We therefore instead first summarize the code source itself by prompting a language model, then cluster based on the similarity of these new descriptions.
Once identified, each tuple is used in two stages. The first is to retrieve relevant already-abstracted functions from the Librarian library: for a given tuple of programs, relevant functions are retrieved by prompting a language model with the entire existing existing library and the original input programs, then instructing it to identify which functions would be useful. The retrieved library functions and the original programs in the tuple are then provided as critical context to the language model to propose a sample budget of $K$ candidate refactorings.
4.2 Rank with compression
Once sampled, all $K$ candidate refactorings are passed through a sort-and-filter evaluation harness to select which one scores the highest on refactor quality and that maintains (or improves) test accuracy compared to the original. If no such candidate exists, the original code is preserved, maintaining existing functionality.
New library functions in the selected refactor are saved into the Librarian library for potential use in downstream refactoring of other programs. We provide the full algorithm in Appendix A.
5 MiniCode
MiniCode evaluates a code agent’s capability to identify abstractions across implementations and design reusable libraries. In order to measure these capabilities, our benchmark presents agents with a collection of code sources, then asks agents to refactor the code sources into a unified library alongside refactorings of the original code sources. There are two key desiderata for collections of code sources: The collections must be compressible, in that there exists a latent shared library abstraction, and verifiable, so that we can measure how well refactored code sources preserve functional correctness. We source problems from two domains: Competition coding and synthesized repositories (Table 1).
Agents are expected to interact with MiniCode via the terminal. We structure the benchmark as refactoring a multi-package Python repository, where each code source in a collection is a Python package in a subdirectory. This requires knowledge of basic bash commands for exploring repositories, editing code, and running tests, as well as how to manage complex, multi-package Python libraries.
CodeContests
Competition problems are crafted with specific variations of algorithmic approaches in mind, resulting in both shared latent concepts and required test cases. As a result, competition coding is naturally both compressible and verifiable.
Each collection consists of multiple code sources, each containing a solution to a competition programming prompt, with associated tests for verification. We take solutions, prompts, and tests from CodeContests [19], a dataset consisting of competition coding problems. Each code source in the collection is structured as a subdirectory consisting of the task description in PROBLEM.md, the initial solution in main.py, and a script to run tests in run.sh. Agents are instructed to create a library.py file, which should is imported into each code source. Since CodeContests has no external dependencies on Python packages, this can be done without explicit structuring as a Python package.
Repositories
For the second domain of synthesized repositories, we propose a data-generating process that first produces project ideas, then generates variations of those project ideas tailored to specialized use-cases. This allows us to control the complexity and degree of overlap between each code source in a collection. Each code source in the repositories domain is comprised of a task description, source code, and test cases for functionality and correctness. MiniCode includes both small repositories (approximately 200 lines of source code each) and large repositories (approximately 6.5k lines of source code each), both of which represent realistic settings where different people with different needs use language models to help them write software for their particular use cases. The refactoring agent is tasked with extracting re-usable functions from across code repositories, and re-writing the original code source repositories to use them.
We approximate the true distribution of code repositories, $p(\rho)$, with a generative process using latent textual library and repository descriptions. The entire space of code repositories is massive. To collect and group code repositories that share meaningful structure that can be abstracted into a useful code library, we (1) sample a textual library description; (2) sample use cases; and last (3) their programmatic implementations. This generative process naturally produces repository collections primed for refactoring. We sample from language models for each step of this process. Prompts are shared in Appendix E.
Each collection consists of multiple synthetic repositories as code sources, where each code source is a subpackage. Collections are obtained by transforming the original repository code sources into a multi-package Python library. Each source’s source code is extracted into a subpackage directory, and its tests into a corresponding test subdirectory. Agents are instructed to write a shared library in a shared subpackage named common. This common shared library must be imported and used to refactor each of the original code sources.
Table 1: MiniCode Statisics
Domain Sources Collections Avg LoC Avg Tests Gen by
Code Contests 300 30 87 10 Humans
Small Repositories 262 22 209 12 o4-mini
Large Repositories 20 10 6,433 101 Claude-Sonnet 3.7
6 Experimental Setup
Grouping Programs into Collections
To facilitate parallel application of Librarian and manage the dataset scale, we assume that semantically distant code sources will have minimal overlap in their optimal library functions. Therefore, our overall approach partitions the dataset into disjoint collections through clustering.
For CodeContests, these collections are constructed from an initial corpus of $\sim$9k problems with Python solutions: We first filter these code sources, removing those whose selected canonical solution is under 10 lines (minimal refactoring potential). For the remaining 4596 solutions we use a language model to generate textual descriptions of canonical solutions—emphasizing reusable components—which are embedded using OpenAI’s text-embedding-ada-002.
Agglomerative Clustering [20] is subsequently applied to these embeddings to partition the code sources into a predefined number of initial clusters, in our case 120. To create uniformly sized experimental units, we subsample each such cluster to form collections of 30 code sources. This collection size was empirically chosen because it balanced between the runtime of Librarian without limiting compression. We select 10 collections that we then use to evaluate our methods.
Code repositories are generated as disjoint collections through the generative process, which first samples project ideas then variations. For small repositories, we up sample $10-15$ code sources then form smaller collections of size $3$-$5$ code sources that exhibit the most intra-cluster similarity according to clustering using the same embedding and clustering technique. For big repositories, we sample two code sources for each of the $10$ collections.
REGAL Baselines
To evaluate the ability of our libraries to support reuse on new problems, we turn to the program synthesis tasks used in REGAL, where learned libraries are added to help the program synthesizer. We evaluate on the two domains published by the authors, Logo and Date. Because our clustering is inspired by REGAL but adds additional complexity, for fair comparison, we keep their setup the same and only augment the training using sample + MDL rerank procedure described in Section 4.1.
Code Contests
To evaluate Librarian on refactoring Code Contests we select 6 collections of 30 code sources (problems). In each collection we group the problems into tuples of size $3$. We set the sample budget to be $K=8$, since our ablations show that with larger $K$ we discover better libraries 3. We use the MDL objective for rankings.
The model used for sampling is OpenAI’s o4-mini [21]. To obtain MDL scores we use Qwen 2.5 7B Instruct [22] as a balance between quality, speed, and cost.
Code Agents
To fairly evaluate performance on the task by state-of-the-art systems, we use coding agents that advertise long-context ability to reason about, write, and refactor code repositories. Specifically, we use Claude Code (Cl) [23] which uses the Claude 3.7 Sonnet model, and OpenAI’s Codex (Cx) [24] which uses o4-mini.
All steps of the initial code sources for small code repositories are done by OpenAI Codex. We empirically observed that Claude Code has an affinity for generating large code repositories– the large code repositories are all generated by Claude Code.
We test whether code agents can refactor collections of code sources autonomously, without human intervention. Refactoring repositories with code agents involves planning and iterative (re-)implementation and testing. Code agents are prompted to perform each of these steps, with feedback from the unit tests. Agents must run and repair unit tests autonomously. We run coding agents multiple times per collection, logging their progress in checklists stored in text files. Our naming convention for agents is “Planner-Executor”. For example, “Cl-Cx” uses the Claude Code agent for planning and OpenAI Codex to implement the plan.
As before we use Qwen 2.5 7B Instruct [22] to obtain MDL scores.
Metric Value
Pass Rate 90.67% ±1.88
Pass Rate Improvement 6.33% ±1.41
MDL Ratio 0.53 ±0.03
Token Ratio 0.66 ±0.04
Library Functions 10.30 ±1.41
Avg Calls per Function 5.17 ±1.08
Percentage Single Use Functions 38.03% ±4.88
Figure 2: Refactoring results for Librarian (w/ $K=8$) averaged over 10 Code Contests collections.
7 Results
In this section, we present the compression and correctness results with Librarian and agent baselines on MiniCode.
MiniCode-CodeContests On CodeContests, Librarian achieves a high final pass rate of 90.67% and significantly improves correctness, with pass rates increasing by 6.33% compared to the original code sources (Table 2). The method yields substantial compression: the refactored code, including the new library, shows an MDL ratio of 0.53 (a 47% reduction in MDL relative to the original). On average, Librarian generates libraries containing approximately 11 functions. These functions demonstrate good reuse, being called by around 5.2 programs on average, although 38.03% of them are used only once within their specific collection context.
Code agents fail to achieve both high correctness and compression on MiniCode-CodeContests. Across collections, the Codex agent achieves an average MDL ratio of 0.83, but a pass rate of 74.16%, much lower than Librarian’s rate of 90.67%. Similarly, the Claude agent reaches a higher pass rate of 82.50%, which is still lower than Librarian, but an MDL ratio of 1.04 which is more complex than the original collection.
MiniCode-Repositories Coding agents results on both small and large repositories are given in Table 2. During experiments, we found that the Claude Code agent provided superior plans for refactoring, while both the Claude Code and Codex agents performed satisfactorily at implementing a plan.
Despite using code agents with a human in the loop, existing code agents fail to produce effective refactorings. Often, the resulting refactored codebase bloats in size and complexity, particularly when using Claude Code for both planning and implementation. Given the poor performance of code agents on these repository-scale refactorings, we do not run the full sample-and-rerank pipeline. Supporting qualitative analysis in Section 8.4 discusses common failure modes of code agents on this task in repository-size refactorings.
Collection Agent MDLPass%
datapipe orig 1.0 1.0
Cl-Cl 9.1 1.0
Cl-Cx 2.2 1.0
state_machine orig 1.0 1.0
Cl-Cl 6.3 0.9
Cl-Cx 1.9 1.0
config_schema orig 1.0 1.0
Cl-Cl 2.5 0
Cl-Cx 2.1 1.0
cli_tools orig 1.0 1.0
Cl-Cl 4.0 1.0
Cl-Cx 3.1 0
cli_form orig 1.0 1.0
Cl-Cl 3.8 1.0
Cl-Cx 2.9 1.0
Collection MDLPass%
filesys_analyzer orig 1.0 1.0
Cl-Cl 1.6 1.0
query_language orig 1.0 1.0
Cl-Cl 1.4 1.0
knowledge_store orig 1.0 1.0
Cl-Cl 1.4 1.0
vm_emulator orig 1.0 1.0
Cl-Cl 1.6 1.0
finance_tracker orig 1.0 1.0
Cl-Cl 1.8 0.6
in_memory_db orig 1.0 1.0
Cl-Cl 1.5 1.0
text_editor orig 1.0 1.0
Cl-Cl 1.2 1.0
Table 2: The correctness (Pass %) and compression (MDL ratio %) of the original and refactored code sources for the agent baselines on a subset of the small and large repository collections in MiniCode. Libraries with syntax or runtime errors receive a pass rate of 0.
8 Analysis
We further analyze parts of Librarian, consider alternative objective functions, analyze the use of clustering while constructing MiniCode-CodeContests, and common failure modes of code agents.
8.1 Can these libraries be reused to solve downstream programming problems?
To understand the practical value of our library extraction work for program synthesis, we augment the existing library learning algorithm, REGAL with sampling and MDL-reranking. In REGAL, the learned library is used to solve downstream holdout program synthesis problems, and a simpler clustering algorithm is used. For fair comparison with REGAL we hold the clustering algorithm constant (using the simpler REGAL approach), but sample $K=5$ refactorings/libraries and rerank by MDL. The resulting system outperforms REGAL on holdout programming problems (Table 3).
These problems however are highly simple, which is representative of classic library learning work: Either drawing simple geometric designs (dataset Logo) or manipulating textual representations of dates (dataset Date). This helped motivate us to consider refactoring whole repositories.
Table 3: Solving downstream program synthesis tasks using learned libraries
Dataset Model Pass Rate
Logo REGAL (gpt-3.5-turbo) 49.3% ±1.1
Librarian (3.5-turbo) 69.9% ±0.9
Date REGAL (gpt-3.5-turbo) 90.2% ±0.5
Librarian (3.5-turbo) 94.7% ±0.7
8.2 What objective function should a library learner use?
We run our method on CodeContests using MDL, number of tokens, and cyclomatic complexity [25] [^2] Cyclomatic complexity (CC) is a classic metric from the software engineering community that analyses the number of paths through a program’s control flow, and unlike MDL is only loosely coupled with program size. as objective functions on 6 collections (Figure 3) While minimizing MDL also minimizes the other two objectives, the converse is not true. This suggests that MDL is a pareto-optimal loss among the three objectives in this experiment.
To confirm that the library does indeed expose shared abstractions, we calculate the average number of times that each library routine is used. Scaling the inference budget to $K=8$ discovers better libraries, reusing each library function on average about 5 times.
In Appendix A.1, we find that MDL is also the objective that best correlates with human preferences in a small-scale human study.
Figure 3: Comparing 3 different objective functions for refactoring (different columns) according to different downstream success metrics (different rows), as a function of refactoring budget (horizontal axes).The values are averaged over 6 collections of CodeContests problems. Row 1: Optimizing perplexity also incidentally optimizes cyclomatic complexity and token count, but that the converse is not true. Row 2: refactored programs pass more test cases, even more than the original code itself. Row 3: increasing the refactoring budget results in more reusable library subroutines (such subroutines are called more times on average). Filtered/Raw: Using/Not Using tests to filter samples.
8.3 Clustering Analysis: CodeContests
We analyze the coherence of the clusters underlying collections in MiniCode-CodeContests. In particular, we compare clustering based on o4-mini generated code source descriptions against task descriptions. Since task descriptions in competition coding problems are designed to hide the algorithmic approach needed to solve problem, we expect that clusters based on code source descriptions are more coherent.
We use two measures to evaluate the thematic coherence of collections: Good collections should group code sources with a (1) concentrated and (2) identifiable set of shared conceptual tags, which for CodeContests are provided as ground truth (trees, graphs, etc.).
Normalized Tag Instance Entropy
This measures the concentration of tag instances within a collection, and is given by the entropy of the tag distribution for a given collection, normalized by the number of distinct tags in that collection. A lower normalized tag instance entropy (closer to 0) indicates higher thematic purity, meaning a small number of tag types are most prevalent.
Herfindahl-Hirschman Index (HHI) for Problem Presence
This measures tag concentration across distinct code sources in a collection. A higher HHI signifies that the problems are collectively characterized by a smaller, more focused set of tags.
We provide the full formal definitions of both measures in Appendix B.
Figure 4 shows our clustering approach yields more thematically coherent clusters, evidenced by achieving lower entropy and higher HHI values across the entire tested range of $N$.
Figure 4: Clustering analysis of 4,596 Code Contest problems, comparing the thematic coherence of clusters formed using our proposed method versus REGAL-style clustering.
8.4 Common Failure Modes of Coding Agents
Failing to follow the plan.
Refactoring large repositories requires editing code over a long time horizon. We instruct agents to plan hierarchically by first recording their plan in natural language, then implementing that plan. However, agents often ignore the plan, particularly with Codex as the implementation agent, potentially refusing to edit the code at all. We hypothesize that this behaviour is caused by the refactoring task itself – the original code source passes all tests, and agents are trained to prioritize correctness over compression.
Failing to use the library.
When the implementation agent does follow the plan, it often fails to do so effectively. Claude Code, for example, will write a library, import it to the proper place, then rewrite the library rather than use it. We provide further examples in Appendix G.
9 Conclusion
We introduce a new benchmark MiniCode and method Librarian for compressing code sources through reusable abstractions. We highlight the challenges that modern models face when producing modular and maintainable code, then present an effective method for using LLMs to do this task in small-scale programs. By framing refactoring as both a design and compression task, our work opens new directions for building more general and scalable code understanding and generation systems. In particular, the structure of MiniCode lends itself well to a reinforcement learning-based task, which synthesizes collections of repositories to refactor then computes rewards based on compression.
Limitations Limitations of this work include the evaluation of synthetic repositories, the poor performance of code agents on refactoring, and the fact that compression may not be correlated with reuse. While our experiments on downstream programming problems partly address the question of reuse, investigating how to measure and encourage reuse for large-scale multi-repo library creation remains an open problem.
Ethics Statement
As with any coding benchmark, our work has the potential to improve automated coding. However, this has the potential to produce malicious code if left to run unchecked. Care should be taken to monitor inputs and outputs to the system.
Acknowledgements
Appendix A Algorithm
Algorithm 1 Refactoring Specialized Programs into a Joint Library
1:Set of independent, specialized programs $P_{initial}=\{\rho_{1},\rho_{2},\dots,\rho_{n}\}$
2:Sample Budget $K$
3:Joint library $\mathcal{L}_{final}$ and set of refactored programs $P_{final}$
4:$\mathcal{C}\leftarrow\text{Cluster}(P_{initial})$
5:$\mathcal{L}_{final}\leftarrow\emptyset$, $P_{final}\leftarrow\emptyset$
6:for all cluster $c\in\mathcal{C}$  do $\triangleright$ Each cluster independently
7:     $T_{C}\leftarrow\text{GroupIntoTuples}(c)$triangleright Get tuples for each cluster
8:     for all tuples $\tau\in T_{C}$ do
9:         $\{f_{retrieved}\}\leftarrow\text{RetrieveRelevantFromLibrary}(\mathcal{L},\tau)$
10:         $S\leftarrow\emptyset$
11:         for $i=1\text{ to }K$ do $\triangleright$ Sample $k$ times
12:              ($(\{f_{new,i}\},\{\rho^{\prime}_{i}\}\})\leftarrow\textsc{Sample}\left(f_{retrieved},\tau\right)$
13:              $S\leftarrow S\cup\{(\{f_{new,i}\},\{\rho^{\prime}_{i}\})\}$
14:         end for
15:         $(f_{best},\{\rho^{\prime}_{best}\})\leftarrow\text{RerankAndSelectBest}(S,\ell\left(\cdot\right))$triangleright Rerank using objective
16:         $\mathcal{L}_{final}\cup\{f_{best}\}$
17:         $P_{final}\cup\{\rho^{\prime}_{best}\}$
18:     end for
19:end for
20:return $\mathcal{L}_{final},P_{final}$
A.1 Objective Functions
Human evaluation best aligns with the MDL objective function.
We collect 19 CodeContests problems and a pair of refactorings for each, and have 4 human judges pick which they prefer. Each refactoring comprises three rewritten programs, plus the library generated by Librarian. Each pair of refactorings passed all testcases, and were chosen to include both the highest and lowest perplexity refactorings. Human judges are authors on this paper, but were ‘blinded’ to not know relative objective function scores for the candidate refactorings.
An MDL objective aligns best with human preferences (Figure 5), although number of tokens is also a reasonable proxy for human judgment. That token count and MDL yield different statistics furthermore implies that the lowest MDL program is not always the shortest program.
Figure 5: Human evaluation of different refactoring objectives. Humans and models give preferences to pairs of refactorings/libraries.
The instruction provided to human evaluators is as follows:
1
2
3## 1. Materials Provided
4
5You will be given a set of files for each example case:
6
7* **‘original_programs.py‘**: This file contains a set of 3 distinct Python programs, each presented with its corresponding problem description/query. This represents the "before" state.
8* **‘v1.py‘**: This file presents the first refactoring approach. It includes:
9 * The 3 refactored versions of the original programs.
10 * A "library" section (e.g., codebank.py or inline) containing helper functions. These helper functions might be retrieved from an existing common library or newly created during this refactoring.
11 * Either the retrieved or the new helper function sections may be _empty_, in case no programs existed in the codebank at the time or if no need helper functions were created by the LLM.
12* **‘v2.py‘**: This file presents the second, alternative refactoring approach. Similar to refactoring_v1.py‘, it includes:
13 * The 3 refactored versions of the original programs (using a different strategy than v1).
14 * A "library" section with its own set of helper functions.
15
16
17**NOTE**: both refactorings had accuracy at least as good as the original programs.
18
19## 2. Your Task
20
21Your primary task is to:
22
231. **Review** the original_programs.py to understand the initial code and the problems being solved.
242. **Analyze** both refactoring_v1.py and refactoring_v2.py‘. Pay close attention to how the original programs have been restructured and what functionalities have been extracted into their respective libraries.
253. **Decide which refactoring (Version 1 or Version 2) you believe is "better,"** based on the evaluation criteria provided below (or your own criteria!).
26
27## 3. Evaluation Criteria: What to Consider for Your Choice
28
29When comparing refactoring_v1.py and refactoring_v2.py‘, please *consider* the following aspects to inform your choice. The "better" refactoring should ideally excel in these areas:
30
31> Most importantly, make sure that the extracted functions are **actually reusable and not too specific.** If the main programs are short, the refactoring is not immediately "better"! Try to think whether the extracted functions could actually be used in a different program down the line.
32
33* **Reusability of Helper Functions :**
34 * **Generality:** Are the new helper functions general-purpose and potentially useful for *other, different* programs and problems beyond the three presented?
35 * **Reuse:** How much were existing helper functions reused?
36 * **Specificity:** Are the functions too specialized to the current set of problems, limiting their broader applicability? _Avoid functions that are essentially just the original program broken out into a "helper."_
37 * Composability
38* **Maintainability:**
39 * Readability & Understandability
40 * Ease of Modification
41 * Separation of Concerns
42
43## 4. What NOT to Focus On:
44
45* **Comments:** Please disregard the presence or absence of comments in the code for this evaluation. These are superficially generated by LLMs in some occasions and could be added manually after with a single pass.
46* **Minor Stylistic Differences:** Do not focus on trivial differences in variable naming or formatting, unless they significantly impact readability or understanding.
47
48## 5. How to Provide Your Feedback
49
50For each example case, please provide:
51
521. **Your Preferred Version:** (e.g., "Version 1" or "Version 2")
Listing 1: Human Evaluation Instruction
Appendix B Collection Coherence Measures
We provide the full definitions of the collection coherence measures.
Normalized Tag Instance Entropy:
This measures the concentration of tag instances within a collection $C$. Let $p_{i}$ be the proportion of the $i$-th unique tag type among all tag instances in $C$, and $D_{C}$ be the number of distinct tag types in $C$. If $D_{C}>1$, the normalized entropy $H_{N}$ is defined as:
$H_{N}=-\frac{\sum_{i=1}^{D_{C}}p_{i}\log_{2}p_{i}}{\log_{2}D_{C}}$ (3)
If $D_{C}\leq 1$, then $H_{N}=0$. Lower $H_{N}$ (closer to 0) indicates higher thematic purity, meaning fewer tag types dominate the bulk of tag mentions.
Herfindahl-Hirschman Index (HHI) for Problem Presence:
This measures tag concentration across distinct problems in a cluster $C$. Let $s_{t}$ be the proportion of problems in $C$ that include tag $t$ (a problem contributes to $s_{t}$ if $t$ is one of its unique tags). A higher HHI signifies that the problems are collectively characterized by a smaller, more focused set of tags.
$\text{HHI}=\sum_{t\in\text{Tags}(C)}s_{t}^{2}$ (4)
where $\text{Tags}(C)$ represents the set of unique tags present in cluster $C$.
Appendix C Benchmark Comparison
We compare our benchmark, MiniCode, to similar benchmarks in Table 4. We define creativity and design as the need to explore diverse solutions in order to find the best solution possible. For example, optimizing for program correctness alone does not require exploring a large solutions space, whereas optimizing a program for speed would. In the case of compressing large code sources, we must explore the large space of shared abstractions afforded by libraries in order to maximize compression.
Table 4: Comparison of Code Benchmarks
Benchmark Creativity/Design Scale
SWE-bench [9] Low Repository
Commit-0 [10] Medium Repository
RefactorBench [11] Low File
ECCO [13] High Function
KernelBench [14] High Function
MiniCode(Ours) High Multi-repository
Appendix D Full MiniCode Repository results
We present the full repository-level results of MiniCode-repository in Tables 5 and 6.
Collection Agent LLoC CC MDL ratio Pass rate
datapipe original 474 116 1.0 1.0
Cl-Cl 1084 341 9.1 1.0
Cl-Cx 645 174 2.2 1.0
state_machine original 619 222 1.0 1.0
Cl-Cl 2271 735 6.3 0.9
Cl-Cx 686 227 1.9 1.0
config_schema original 949 284 1.0 1.0
Cl-Cl 922 339 2.5 failed
Cl-Cx 626 207 2.1 1.0
cli_tools original 875 300 1.0 1.0
Cl-Cl 2925 909 4.0 1.0
Cl-Cx 5848 2378 3.1 failed
cli_form original 352 174 1.0 1.0
Cl-Cl 855 342 3.8 1.0
Cl-Cx 1366 548 2.9 1.0
Table 5: Full results on MiniCode-repositories small.
Collection LLoC CC MDL ratio Pass rate
file_system_analyzer original 3911 1338 1.0 1.0
refactored 6495 2218 1.6 1.0
query_language_interpreter original 5181 2105 1.0 1.0
refactored 7193 2966 1.4 1.0
personal_knowledge_management original 5341 1832 1.0 1.0
refactored 6357 2364 1.4 1.0
virtual_machine_emulator original 6841 2283 1.0 1.0
refactored 10108 3580 1.6 1.0
personal_finance_tracker original 4306 1482 1.0 1.0
refactored 7443 2512 1.8 0.6
in_memory_database original 4565 1671 1.0 1.0
refactored 6667 2556 1.5 1.0
command_line_task_manager original 7964 2921 1.0 1.0
refactored 10387 3899 1.3 1.0
text_editor original 4324 1323 1.0 1.0
refactored 5792 1822 1.2 1.0
concurrent_task_scheduler original 10350 3935 1.0 1.0
refactored 18653 7165 1.9 0.8
incremental_backup_system original 4587 1482 1.0 1.0
refactored 5365 1788 1.3 failed
Table 6: Full results on MiniCode-repositories large.
Appendix E Data Generating Prompts
Prompt to generate general library text descriptions.
1I need ideas for Python libraries that can be implemented by language models. These libraries should:
2
31. Be implementable using only Pythons standard library - no external dependencies
42. Have enough complexity to demonstrate sophisticated code design (10-20 functions/methods)
53. Include room for interpretation, so that different implementations can be unique while sharing core functionality
64. Have clear, practical utility that solves a real programming need
75. Be realistically implementable by an intelligent language model
86. Be testable with pytest
97. Include opportunities for different design approaches (functional vs OOP, etc.)
10
11For each library, provide a description that outlines:
12- The problem domain and core purpose
13- Key required functionality (without being too prescriptive about implementation details)
14- Potential use cases that demonstrate practical applications
15- Suggested extension points where implementers could add their creative spin
16
17Please generate several proposals in markdown, following this format:
18
19‘‘‘file:<library_name>/DESCRIPTION.md
20# <Library Name>
21
22## Purpose and Motivation
23<3-5 sentences on what problem this library solves and why its useful>
24
25## Core Functionality
26<Description of 4-6 high-level key features/capabilities without specifying exact implementation>
27‘‘‘
28
29Be creative! Focus on domains where standard Python libraries provide enough building blocks but where a well-designed abstraction layer would add significant value.
Listing 2: Prompt to generate library descriptions
1
2Consider a code repository designed to support the following task description:
3‘‘‘
4{task_content}
5‘‘‘
6
7Please list a couple dozen features that would be useful for the repository. Include the specified features as well as several others.
8
9List the suggested feature names and descriptions in the following format:
10
111: <feature_1_name>: <feature_1_description>
122: <feature_2_name>: <feature_2_description>
133: <feature_3_name>: <feature_3_description>
14...
1530: <feature_30_name>: <feature_30_description>
16"""
17
18persona_prompt_template = """Consider the following features of a code repository:
19{listed_features}
20
21Think about several possibilities for what kind of person might use this code repository and what they might use it for. Please write several brief descriptions for the code repository in first person, formatted in markdown as follows:
22‘‘‘file:{library_name}/<persona_name>/TASK.md
23# The Task
24
25I am a <...> I want to be able to <...> This code repository <...>
26
27# The Requirements
28
29* ‘<function_name>‘ : <feature description>
30* ...
31‘‘‘
32
33Be creative! Write the task description in the style of the proposed persona. Be as exhaustive as possible in including the listed features in the task descriptions requirements.
Listing 3: Prompt to generate potential uses and features given a library description
1
2
3I need you to implement a Python solution and COMPREHENSIVE suite of tests based on the following task files.
4Your code must pass the tests provided.
5
6{task_content}
7
8CRITICAL FORMATTING INSTRUCTIONS:
91. You MUST format ALL code files exactly as shown below - no exceptions
102. Start each file with the markdown codeblock marker, followed by "file:" and the relative path
113. End each file with the closing markdown codeblock marker
124. Do not use any other format or markdown variations
135. For test files, do not put them in a subdirectory-- keep them in the outermost level.
14
15For each source code file:
16
17‘‘‘file:<relative_file_path>
18<file_content>
19‘‘‘
20
21For each test file:
22
23‘‘‘file:<relative_file_path starting with test_>
24<test_file_content>
25‘‘‘
26
27IMPORTANT:
28- The opening format must be exactly: ‘‘‘file:path/to/file.py
29- Do not add language indicators like ‘‘‘python
30- Do not add explanations between files
31- Each file must be contained within its own codeblock with the precise format shown above
32- The system parsing your response requires this exact format to function properly
33
34EXAMPLE OUTPUT FORMAT:
35‘‘‘file:mymodule/mymodule.py
36def example_function():
37 return "This is a sample function"
38‘‘‘
39
40‘‘‘file:test_utils.py
41import pytest
42from mymodule.mymodule import example_function
43
44def test_example_function():
45 assert example_function() == "This is a sample function"
46‘‘‘
47
48Begin your implementation now, following these formatting rules precisely.
Listing 4: Prompt to generate initial implementation
1
2I need you to fix the implementation of the following code that is failing tests.
3
4# Current Implementation:
5{src_code_content}
6
7# Test Files:
8{test_content}
9
10# Failed Tests:
11{failed_test_details}
12
13# Test Output (if available):
14{test_output}
15
16Please carefully analyze the errors and test failures. Pay special attention to:
171. The exact assertion failures or error messages
182. What the tests expect vs. what your current implementation provides
193. Any edge cases or special conditions you might have missed
20
21Your task is to fix the implementation to make all tests pass. For each file that needs to be modified,
22provide the content in the following format:
23
24‘‘‘file:<relative_file_path>
25<file_content>
26‘‘‘
27
28Where <relative_file_path> is the relative path to the file and <file_content> is the updated content of the file.
29Focus on fixing the specific issues identified in the errors and failed tests while maintaining the overall structure of the code.
30
31IMPORTANT: Make targeted changes to address the specific failing test cases. Make sure your implementation passes all test cases,
32including any edge cases or special conditions mentioned in the tests. Be sure to output code in the specified format.
Listing 5: Prompt to fix code implementation, given pytest output.
Appendix F Refactoring examples of Librarian on Code Contests
F.1 Example 1
In code snippets LABEL:code:v1, LABEL:helper:v1, LABEL:code:v2, LABEL:helper:v2 one example of 2 refactoring versions. Specifically, the versions are both passing at least as many test cases as the original and they have the biggest difference in MDL among all the sample refactorings for that tuple. Sample + rerank filtering selected refactoring V2. You can observe that refactoring V1 introduces some problem specific functions like build_max_beauty_perm(), while refactoring V2 sticks to more generally useful functions.
1# ==== NEW HELPER FUNCTIONS ====
2def compute_full_mask(i):
3 """Return mask of all 1s of the bit-length of i."""
4 return (1 << i.bit_length()) - 1
5
6def build_max_beauty_perm(n):
7 """Build permutation of 0..n maximizing sum of i^p[i]."""
8 ans = [0] * (n + 1)
9 used = set()
10 for i in range(n, -1, -1):
11 if i in used:
12 continue
13 mask = compute_full_mask(i)
14 j = i ^ mask
15 ans[i], ans[j] = j, i
16 used.add(i)
17 used.add(j)
18 beauty = sum(i ^ ans[i] for i in range(n + 1))
19 return ans, beauty
20
21def solve_xor_sum(u, v):
22 """
23 Find shortest array whose xor is u and sum is v.
24 Return list or None if impossible.
25 """
26 if u > v or (v - u) % 2:
27 return None
28 if u == v:
29 return [] if u == 0 else [u]
30 x = (v - u) // 2
31 # try two elements
32 if ((u + x) ^ x) == u:
33 return [u + x, x]
34 # fallback to three elements
35 return [u, x, x]
36
37def build_trie(keys):
38 """
39 Build a binary trie with counts for 30-bit numbers.
40 Each node: [left_index, right_index, count].
41 """
42 tree = [[0, 0, 0]]
43 for x in keys:
44 now = 0
45 tree[now][2] += 1
46 for i in range(29, -1, -1):
47 b = (x >> i) & 1
48 if tree[now][b] == 0:
49 tree[now][b] = len(tree)
50 tree.append([0, 0, 0])
51 now = tree[now][b]
52 tree[now][2] += 1
53 return tree
54
55def trie_pop_min_xor(tree, x):
56 """
57 Pop one key from trie to minimize x^key and return that minimal xor.
58 Decrements counts along the path.
59 """
60 now = 0
61 res = 0
62 for i in range(29, -1, -1):
63 b = (x >> i) & 1
64 nxt = tree[now][b]
65 if nxt and tree[nxt][2] > 0:
66 now = nxt
67 else:
68 now = tree[now][b ^ 1]
69 res |= (1 << i)
70 tree[now][2] -= 1
71 return res
Listing 6: Version 1, New Helpers
1
2# ########## PROGRAM: node_16:cc_python_16 ##########
3
4from codebank import *
5
6def main():
7 import sys
8 data = sys.stdin.readline()
9 if not data:
10 return
11 n = int(data)
12 perm, beauty = build_max_beauty_perm(n)
13 print(beauty)
14 print(*perm)
15
16if __name__ == "__main__":
17 main()
18
19# ########## PROGRAM: node_19:cc_python_19 ##########
20
21from codebank import *
22
23def main():
24 import sys
25 data = sys.stdin.readline
26 n = int(data())
27 A = list(map(int, data().split()))
28 P = list(map(int, data().split()))
29 trie = build_trie(P)
30 O = [trie_pop_min_xor(trie, a) for a in A]
31 print(*O)
32
33if __name__ == "__main__":
34 main()
35
36# ########## PROGRAM: node_25:cc_python_25 ##########
37
38from codebank import *
39
40def main():
41 import sys
42 u, v = map(int, sys.stdin.readline().split())
43 res = solve_xor_sum(u, v)
44 if res is None:
45 print(-1)
46 else:
47 print(len(res))
48 if res:
49 print(*res)
50
51if __name__ == "__main__":
52 main()
Listing 7: Version 1, Refactored Programs
1# ==== NEW HELPER FUNCTIONS ====
2def compute_complement(i):
3 return i ^ ((1 << i.bit_length()) - 1)
4
5def trie_add(trie, x, max_bit):
6 trie[0][2] += 1
7 now = 0
8 for i in range(max_bit, -1, -1):
9 bit = (x >> i) & 1
10 if trie[now][bit] == 0:
11 trie[now][bit] = len(trie)
12 trie.append([0, 0, 0])
13 now = trie[now][bit]
14 trie[now][2] += 1
15
16def trie_find_min_xor(trie, x, max_bit):
17 now = 0
18 ans = 0
19 for i in range(max_bit, -1, -1):
20 bit = (x >> i) & 1
21 if trie[now][bit] and trie[trie[now][bit]][2] > 0:
22 now = trie[now][bit]
23 else:
24 now = trie[now][bit ^ 1]
25 ans |= (1 << i)
26 trie[now][2] -= 1
27 return ans
Listing 8: Version 2, New Helpers
1# ########## PROGRAM: node_16:cc_python_16 ##########
2
3from codebank import *
4
5def main():
6 import sys
7 input = sys.stdin.readline
8 n = int(input())
9 ans = [-1] * (n + 1)
10 for i in range(n, -1, -1):
11 if ans[i] == -1:
12 z = compute_complement(i)
13 ans[i] = z
14 ans[z] = i
15 m = sum(i ^ ans[i] for i in range(n + 1))
16 print(m)
17 print(*ans)
18
19if __name__ == "__main__":
20 main()
21
22# ########## PROGRAM: node_19:cc_python_19 ##########
23
24from codebank import *
25
26def main():
27 import sys
28 input = sys.stdin.readline
29 n = int(input())
30 A = list(map(int, input().split()))
31 P = list(map(int, input().split()))
32 max_bit = max(max(A, default=0), max(P, default=0)).bit_length() - 1
33 trie = [[0, 0, 0]]
34 for x in P:
35 trie_add(trie, x, max_bit)
36 res = [trie_find_min_xor(trie, x, max_bit) for x in A]
37 print(*res)
38
39if __name__ == "__main__":
40 main()
41
42# ########## PROGRAM: node_25:cc_python_25 ##########
43
44from codebank import *
45
46def main():
47 u, v = map(int, input().split())
48 if u > v or ((v - u) & 1):
49 print(-1)
50 elif u == 0 and v == 0:
51 print(0)
52 elif u == v:
53 print(1)
54 print(u)
55 else:
56 w = (v - u) // 2
57 if (w & u) == 0:
58 d = u + w
59 print(2)
60 print(d, w)
61 else:
62 print(3)
63 print(u, w, w)
64
65if __name__ == "__main__":
66 main()
Listing 9: Version 2, Refactored Programs
F.2 Example 2
In code snippets LABEL:code:v1_2, LABEL:helper:v1_2, LABEL:code:v2_2, LABEL:helper:v2_2 is another example of 2 refactorings where V1 was better according to Librarian. We can observe that V2 creates helper functions that are overly specific to the problem. You can see that refactoring V2 introduces overly specialized functions like dijkstra_special() or compute_min_moves_opposite_parity(). In comparison, refactoring V1 generates only general versions of these functions (e.g. dijkstra()).
1
2# ==== NEW HELPER FUNCTIONS ====
3def read_ints():
4 return list(map(int, input().split()))
5
6def build_adj_undirected(n, edges):
7 adj = [[] for _ in range(n)]
8 for u, v, w in edges:
9 adj[u].append((v, w))
10 adj[v].append((u, w))
11 return adj
12
13def dijkstra(adj, src):
14 from heapq import heappush, heappop
15 INF = 10**18
16 n = len(adj)
17 dist = [INF]*n
18 parent = [-1]*n
19 dist[src] = 0
20 heap = [(0, src)]
21 while heap:
22 d, u = heappop(heap)
23 if d > dist[u]:
24 continue
25 for v, w in adj[u]:
26 nd = d + w
27 if nd < dist[v]:
28 dist[v] = nd
29 parent[v] = u
30 heappush(heap, (nd, v))
31 return dist, parent
32
33def reconstruct_path(parent, dest):
34 path = []
35 u = dest
36 while u != -1:
37 path.append(u+1)
38 u = parent[u]
39 return path[::-1]
40
41def multi_source_bfs(neighbors, sources):
42 from collections import deque
43 n = len(neighbors)
44 dist = [-1]*n
45 dq = deque()
46 for u in sources:
47 if dist[u] == -1:
48 dist[u] = 0
49 dq.append(u)
50 while dq:
51 u = dq.popleft()
52 for v in neighbors[u]:
53 if dist[v] == -1:
54 dist[v] = dist[u] + 1
55 dq.append(v)
56 return dist
Listing 10: Version 1, New Helpers
1# ########## PROGRAM: node_16:cc_python_16 ##########
2
3from codebank import *
4
5def main():
6 import heapq
7 n, m = read_ints()
8 edges = [(u-1, v-1, w) for u, v, w in (read_ints() for _ in range(m))]
9 adj = build_adj_undirected(n, edges)
10 INF = 10**20
11 dist = [INF]*n
12 dist[0] = 0
13 last_w = [0]*n
14 heap = [(0, 0)]
15 while heap:
16 d, u = heapq.heappop(heap)
17 if d > dist[u]:
18 continue
19 # record last edges
20 for v, w in adj[u]:
21 last_w[v] = w
22 # expand two-edge moves
23 for v, w1 in adj[u]:
24 tw = last_w[v]
25 for x, w2 in adj[v]:
26 nd = d + (tw + w2)**2
27 if nd < dist[x]:
28 dist[x] = nd
29 heapq.heappush(heap, (nd, x))
30 out = []
31 for x in dist:
32 out.append(str(x if x < INF else -1))
33 print(" ".join(out))
34
35if __name__ == "__main__":
36 main()
37
38# ########## PROGRAM: node_17:cc_python_17 ##########
39
40from codebank import *
41
42def main():
43 n, m = read_ints()
44 edges = [(u-1, v-1, w) for u, v, w in (read_ints() for _ in range(m))]
45 adj = build_adj_undirected(n, edges)
46 dist, parent = dijkstra(adj, 0)
47 if dist[n-1] >= 10**18:
48 print(-1)
49 else:
50 path = reconstruct_path(parent, n-1)
51 print(*path)
52
53if __name__ == "__main__":
54 main()
55
56# ########## PROGRAM: node_19:cc_python_19 ##########
57
58from codebank import *
59
60def main():
61 n = int(input())
62 a = read_ints()
63 # build reversed graph: for each move i->j, add edge j->i
64 neighbors = [[] for _ in range(n)]
65 for i, val in enumerate(a):
66 for j in (i - val, i + val):
67 if 0 <= j < n:
68 neighbors[j].append(i)
69 # BFS from all even and all odd positions separately
70 even_sources = [i for i, val in enumerate(a) if val % 2 == 0]
71 odd_sources = [i for i, val in enumerate(a) if val % 2 == 1]
72 dist_even = multi_source_bfs(neighbors, even_sources)
73 dist_odd = multi_source_bfs(neighbors, odd_sources)
74 # for odd a[i], answer is dist to nearest even => dist_even; else dist_odd
75 ans = [dist_even[i] if a[i] % 2 == 1 else dist_odd[i] for i in range(n)]
76 print(*ans)
77
78if __name__ == "__main__":
79 main()
Listing 11: Version 1, Refactored Programs
1#
2# ==== NEW HELPER FUNCTIONS ====
3def read_ints():
4 return list(map(int, input().split()))
5
6def build_undirected_weighted_graph(n, m):
7 from collections import defaultdict
8 adj = defaultdict(list)
9 for _ in range(m):
10 u, v, w = read_ints()
11 u -= 1; v -= 1
12 adj[u].append((v, w))
13 adj[v].append((u, w))
14 return adj
15
16def dijkstra(adj, src, n):
17 import heapq
18 INF = 10**18
19 dist = [INF]*n
20 parent = [-1]*n
21 visited = [False]*n
22 dist[src] = 0
23 heap = [(0, src)]
24 while heap:
25 d, u = heapq.heappop(heap)
26 if visited[u]:
27 continue
28 visited[u] = True
29 for v, w in adj.get(u, ()):
30 nd = d + w
31 if nd < dist[v]:
32 dist[v] = nd
33 parent[v] = u
34 heapq.heappush(heap, (nd, v))
35 return dist, parent
36
37def reconstruct_path(parent, dest):
38 path = []
39 while dest != -1:
40 path.append(dest+1)
41 dest = parent[dest]
42 return path[::-1]
43
44def dijkstra_special(e, n, src):
45 import heapq
46 INF = 10**18
47 d = [INF]*n
48 d[src] = 0
49 heap = [(0, src)]
50 while heap:
51 cd, v = heapq.heappop(heap)
52 if cd > d[v]:
53 continue
54 td = {}
55 for u, w in e.get(v, ()):
56 td[u] = w
57 for u, w1 in td.items():
58 for x, w2 in e.get(u, ()):
59 cost = cd + (w1 + w2)**2
60 if cost < d[x]:
61 d[x] = cost
62 heapq.heappush(heap, (cost, x))
63 return d
64
65def compute_min_moves_opposite_parity(a):
66 from collections import deque
67 n = len(a)
68 go = [[] for _ in range(n)]
69 ans = [-1]*n
70 q = deque()
71 for i, val in enumerate(a):
72 for j in (i - val, i + val):
73 if 0 <= j < n:
74 if (a[j] % 2) != (val % 2):
75 ans[i] = 1
76 q.append(i)
77 break
78 else:
79 go[j].append(i)
80 while q:
81 u = q.popleft()
82 for v in go[u]:
83 if ans[v] == -1:
84 ans[v] = ans[u] + 1
85 q.append(v)
86 return ans
Listing 12: Version 2, New Helpers
1#
2# ########## PROGRAM: node_16:cc_python_16 ##########
3
4from codebank import *
5
6def main():
7 n, m = read_ints()
8 e = {}
9 for _ in range(m):
10 u, v, w = read_ints()
11 u -= 1; v -= 1
12 e.setdefault(u, []).append((v, w))
13 e.setdefault(v, []).append((u, w))
14 d = dijkstra_special(e, n, 0)
15 print(" ".join(str(-1 if x >= 10**18 else int(x)) for x in d))
16
17if __name__ == "__main__":
18 main()
19
20# ########## PROGRAM: node_17:cc_python_17 ##########
21
22from codebank import *
23
24def main():
25 n, m = read_ints()
26 adj = build_undirected_weighted_graph(n, m)
27 dist, parent = dijkstra(adj, 0, n)
28 if dist[n-1] >= 10**18:
29 print(-1)
30 else:
31 path = reconstruct_path(parent, n-1)
32 print(" ".join(map(str, path)))
33
34if __name__ == "__main__":
35 main()
36
37# ########## PROGRAM: node_19:cc_python_19 ##########
38
39from codebank import *
40
41def main():
42 n = int(input())
43 a = read_ints()
44 ans = compute_min_moves_opposite_parity(a)
45 print(" ".join(map(str, ans)))
46
47if __name__ == "__main__":
48 main()
Listing 13: Version 2, Refactored Programs
Appendix G Code Agent Failure Modes
1
2from datapipe.core import (
3 tumbling_window,
4 sliding_window as _sliding_window,
5 add_serializer,
6 throttle_upstream as _throttle_upstream,
7 watermark_event_time as _watermark_event_time,
8 ...
9)
10...
11
12def throttle_upstream(max_size):
13 """
14 Apply backpressure to slow data ingestion if downstream stages are overloaded.
15
16 Args:
17 max_size: maximum queue size or rate limit
18
19 Returns:
20 Decorator function
21 """
22 def decorator(func):
23 from functools import wraps
24
25 @wraps(func)
26 def queue_wrapper(q, *args, **kwargs):
27 try:
28 size = q.qsize()
29 if size > max_size:
30 import time
31 time.sleep(0.01)
32 except Exception:
33 pass
34 return func(q, *args, **kwargs)
35
36 return queue_wrapper
37
38 return decorator
39
40def watermark_event_time(events, allowed_lateness):
41 """
42 Assign event-time watermarks to handle late data correctly.
43
44 Args:
45 events: list of dicts with timestamp
46 allowed_lateness: seconds of allowed lateness
47
48 Returns:
49 Events with watermark annotations
50 """
51 # Ensure we return the appropriate format with is_late field
52 result = []
53 for e in events:
54 tagged = dict(e)
55 max_ts = max(ev[’timestamp’] for ev in events)
56 watermark = max_ts - allowed_lateness
57 tagged[’watermark’] = watermark
58 tagged[’is_late’] = e[’timestamp’] < watermark
59 result.append(tagged)
60 return result
61...
Listing 14: Claude fails to use imports and instead re-implements the function.

### v3

@@
 Maintainable and general software allows developers to build robust applications efficiently, yet achieving these qualities often requires refactoring specialized solutions into reusable components.
-This challenge becomes particularly relevant as code agents become increasingly accurate at solving isolated programming problems.
-We investigate code agents’ capacity to refactor code in ways supporting growth and reusability.
-We present both a method and a benchmark for refactoring: Librarian, a sample-and-rerank method for generating reusable libraries, and MiniCode, a benchmark where code agents must minimize and refactor multiple independent solutions into a joint library.
-Compared to state-of-the-art code agents, Librarian achieves strong results on both compression and correctness on MiniCode, obtaining compression rates 1.6-2x better than coding agents while also improving correctness.
-We open-source our code and benchmark at https://code-refactor.github.io/.
+This challenge becomes particularly relevant as code agents become used to solve isolated one-off programming problems.
+We investigate code agents’ capacity to refactor code in ways that support growth and reusability.
+We first investigate what makes a good refactoring, finding via simulation results and a human study that Minimum Description Length best correlates with preferable refactorings.
+We then present both a benchmark and a method for refactoring: MiniCode, a benchmark where multiple files must be refactored into a shared library, and Librarian, a sample-and-rerank method for generating reusable libraries.
+We compare Librarian to state-of-the-art library generation methods, and study it on real-world code bases.
 1 Introduction
 Writing code is mainly a matter of rewriting code: debugging, refactoring, optimizing, and other activities within the software engineering lifecycle.
-But poor rewrites incur technical debt, with such debt costing up to \$2 trillion annually [1].
+But poor rewrites incur technical debt, with such debt costing up to \$2 trillion annually (Tews, ).
 This problem will likely worsen as language models become increasingly responsible for generating code, because they excel at solving isolated programming problems, but their context length demands a myopic view of the codebase.
-It is therefore valuable to understand not just the ability of language models to solve programming problems, but also their ability to rewrite and refactor code in ways that support growth and reuse.
-Effective code refactoring at scale is a design problem.
-When refactoring codebases, developers must navigate design decisions around concerns such as generality, re-usability, and maintainability.
+Effective code refactoring at scale is a design problem whose concerns center on re-usability and maintainability.
 A classic example illustrates this design challenge: Human programmers often create overly-specialized, redundant solutions to similar problems and would benefit from redesigning specialized solutions into a shared library.
-This consolidation requires careful design decisions about the right level of abstraction — neither too specific nor too general — and appropriate interfaces that balance flexibility with usability.
-Here we focus on refactoring multiple code sources into a reusable software library, and pose the following question: To what extent can code agents address this problem, both within human-written codebases, and also in language model-generated code?
-To answer that question, we develop a new method and a benchmark.
-This goes beyond past work [2, 3, 4, 5, 6, 7, 8] in library learning that synthesized subroutines across small programs in i.e.
-$\lambda$-calculus, instead tackling the more naturalistic problem of redesigning large bodies of code written in contemporary high-level languages, such as Python, producing classes, methods, and helper functions in the style of a human-written library.
-We propose a simple method, Librarian (Figure 1), which samples possible code rewrites then reranks those samples based on criteria designed to capture what it means to have a good refactoring.
-To generate potential rewrites, we develop methods for clustering pieces of code together that share common structure, so that a succinct prompt can rewrite them jointly into their refactored form.
-To evaluate our method and systematically assess the capability of current systems to perform such design-intensive refactorings, we introduce a new benchmark, MiniCode.
-Existing benchmarks such as SWE-Bench [9], Commit0 [10], and RefactorBench [11] primarily focus on functional correctness and do not directly evaluate an agent’s ability to perform large-scale code consolidation or library design.
-MiniCode fills this gap by explicitly challenging code agents to minimize and refactor multiple specialized code sources into a unified library.
-MiniCode requires agents to design libraries that expose a general interface shared across multiple use cases in two domains: competition coding and repository-level Python applications.
-This involves optimizing not only for functional correctness but also for broader software engineering objectives such as reusability and maintainability, making the task substantially more open-ended than prior benchmarks.
-Our results show that state-of-the-art code agents, based on Claude 3.7 Sonnet and o4-mini, struggle to jointly preserve correctness and improve reusability across both domains of MiniCode.
-In the competition coding domain, our method Librarian improves refactoring quality by 1.89x while also enhancing correctness.
-However, on the repository-level refactoring, even the strongest agents fail to produce high-quality refactorings, highlighting a substantial gap between current capabilities and the demands of design-oriented code rewriting.
-Addressing this challenge remains an open and important direction for future research.
-Figure 1: Overview of the problem that we study and the general structure of its solutions.
-Given a collection of different code sources, where a source is either program or repository, we design a modular and reusable library.
-To do this we cluster together related programs into tuples (left), sample different rewrites using a language model (right), and select the rewrites which optimize various criteria, such as compression, while validating candidate rewrites using test cases (pass rate).
+Here we focus on refactoring multiple files into a reusable software library, which raises two questions: (1) How should we quantify the quality of a library refactoring, and (2) How can we use LLMs to refactor specialized code into reusable libraries?
+To answer those question, we develop a new method and a benchmark.
+This goes beyond past work in library learning (Wong et al., 2021; Stengel-Eskin et al., 2024; Ellis et al., 2021; Bowers et al., 2023; Dechter et al., 2013; Grand et al., 2023; Liang et al., 2010b) which synthesizes subroutines across small programs in i.e.
+$\lambda$-calculus, because instead we address the more naturalistic problem of redesigning code written in contemporary high-level languages, such as Python, producing classes, methods, and helper functions in the style of a human-written library.
+We develop a method, Librarian (Figure˜1), which samples possible code rewrites and then reranks those samples based on criteria designed to capture a good refactoring.
+To generate potential rewrites, we develop methods for clustering pieces of code together that share common structure so that a succinct prompt can rewrite them jointly into their refactored form.
+To find strong criteria for ranking potential rewrites, we study a variety of metrics across machine learning and software engineering, both on programming benchmarks and via a human study.
+To evaluate our method and systematically assess the capability of current agents to generate libraries, we introduce a new benchmark, MiniCode, which addresses three key desiderata missing from existing benchmarks.
+First, open-ended design: unlike SWE-Bench (Jimenez et al., 2024), Commit0 (Zhao et al., 2025), and RefactorBench (Gautam et al., 2025) which primarily focus on functional correctness, MiniCode presents an unconstrained library design problem.
+Agents create a library that can be imported back into a repository, with complete freedom to design the interface and implementation from scratch—optimizing for software engineering objectives like reusability and maintainability.
+Second, verifiability: we ensure objective evaluation by retaining the unit tests from all repositories that will import the designed library, allowing us to verify that the solutions work correctly across multiple use cases.
+Third, large context: agents must understand and synthesize information from multiple code sources (files) simultaneously to design a unified library that consolidates specialized code sources into a general interface.
+Figure 1: Overview of the refactoring problem.
+A refactoring task comprises a set of files.
+We refactor the files by designing a new library.
+Candidate refactorings are evaluated based on a refactoring metric, and are expected to maintain correctness of the original code sources (pass rate).
+We explore several refactoring metrics in this paper.
+We contribute the following:
+  1. 1.
+    Study of different metrics for library creation via a user study and program synthesis benchmarks, with the surprising finding that neural language models offer a stronger signal of refactoring quality than classic metrics from the software engineering community.
+  2. 2.
+    A new benchmark, MiniCode, covering both competition programming and real-world machine learning programs from the Transformers and Diffusers libraries.
+  3. 3.
+    A new algorithm, Librarian, which outperforms prior library learners on existing benchmarks and scales to complex codebases.
+  4. 4.
+    Refactorings of real-world codebases widely used in the machine learning community.
+Our method successfully refactors the Transformers and Diffusers libraries from Huggingface to be shorter and more reusable.
+To the best of our knowledge this is the first time library learning has have been successfully applied to real-world software projects.
 2 Related work
-Repo-level language model coding agents.
-Recent work has explored the application of language models to repository-level software engineering tasks.
-Existing benchmarks include SWE-bench [9], which evaluates models on their ability to resolve real-life GitHub issues, and Commit-0 [10], which challenges AI systems to reconstruct entire code repositories from scratch.
-Such benchmarks primarily evaluate functional correctness via unit tests, without assessing the quality or maintainability of the resulting codebase.
-RefactorBench [11] takes a step in this direction by benchmarking the ability to follow specific refactoring instructions.
-Our work differs in that we expect models to perform a more open-ended task: Redesigning code to be more modular and compact by discovering and drawing out reused abstractions.
 Library Learning.
 Systems which perform library learning research discover shared abstractions across a large number of small programs, which they use to automatically define new subroutines.
-Systems such as DreamCoder [4], Trove [12], LiLo [7], and REGAL [3] automatically construct such libraries with the goal of making future program synthesis tasks easier to solve, once the learned library is in hand.
-Our work is closest to REGAL [3], which clusters related code and refactors using language models.
+Systems such as DreamCoder (Ellis et al., 2021), Trove (Wang et al., 2024), LiLo (Grand et al., 2023), and REGAL (Stengel-Eskin et al., 2024) automatically construct such libraries with the goal of making future program synthesis tasks easier to solve, once the learned library is in hand.
+Our work is closest to REGAL (Stengel-Eskin et al., 2024), which clusters related code and refactors using language models.
 However, existing library learning approaches have primarily been demonstrated in small-scale, constrained domains, limiting their applicability to typical software engineering tasks, such as consolidating multiple repositories into cohesive libraries.
-By framing library learning within the context of realistic, large-scale code repository development, we expand the relevance of library learning to everyday software engineering practice.
+By framing library learning within the context of realistic software development, we expand the relevance of library learning to the everyday practice of software engineering.
+Repo-level coding benchmarks.
+Recent work has explored the application of language models to repository-level software engineering tasks.
+Existing benchmarks include SWE-bench (Jimenez et al., 2024), which evaluates models on their ability to resolve real-life GitHub issues, and Commit-0 (Zhao et al., 2025), which requires agents to fill in function definitions.
+Such benchmarks primarily evaluate functional correctness via unit tests, without assessing the quality or maintainability of the resulting codebase.
+RefactorBench (Gautam et al., 2025) takes a step in this direction by benchmarking the ability to follow specific refactoring instructions.
+Our work differs by requiring models to perform a more open-ended task: Redesigning code to be more modular and compact by discovering reused abstractions, while retaining verifiability by re-using downstream unit tests.
+Additionally, libraries must be created without any scaffolding limitations such as preexisting function definitions affording more design freedom than Commit-0.
 Program optimization.
-While our goal is to optimize the quality of libraries, other works focus on improving code speed through correctness-preserving transformations [13, 14, 15].
-Both forms of program optimization, compression and speed, require more exploration than only optimizing for correctness, as there does not exist a ground-truth answer.
+While our goal is to optimize the quality of libraries, other works focus on improving execution speed through correctness-preserving transformations (Waghjale et al., 2024; Ouyang et al., 2025; Schkufza et al., 2013).
+Both forms of program optimization, compression and speed, are more open-ended than optimizing only for correctness, as there does not exist a ground-truth answer.
 Prior work on program optimization benchmarks study code at the file level.
-We propose a benchmark that transforms programs at a larger scale, across multiple code repositories.
+We propose a benchmark that transforms programs at a larger scale, across multiple code files.
 3 Problem Statement
-In this section, we propose a refactoring task: Given multiple code sources that contain problem-specific implementations, the goal is to create a cohesive library that captures shared abstractions.
-This library must reduce the total code size while supporting all original use cases, potentially opening up new use cases as well by mining and formalizing latent shared abstractions.
-This is accomplished by searching for refactorings that are both correct and simple.
-Correctness is straightforward to define as the fraction of unit tests passed, but simplicity is more elusive.
-One potential measure of simplicity is counting the total number of tokens in the proposed library and refactored code.
-However, just minimizing program size has obvious failure modes: code should also be natural, elegant, and extensible, which can be in tension with merely finding the shortest program.
+Given $N$ related files $\left\{\rho_{n}\right\}_{n=1}^{N}$, the goal is to create a library $\mathcal{L}$ that captures shared abstractions.
+The original files $\left\{\rho_{n}\right\}_{n=1}^{N}$, which define the refactoring problem, we call a task.
+The new library must support all original use cases in the task by extracting latent shared abstractions.
+This is accomplished by searching for refactorings that are both correct and ‘natural’.
+Correctness is straightforward to define via unit tests, but naturalness is more challenging to quantify.
+Shorter code is potentially simpler and less redundant.
+One potential metric is to count the number of tokens, lines-of-code, or syntax tree nodes in the proposed library and refactored code  (Dechter et al., 2013; Polozov & Gulwani, 2015; Bowers et al., 2023; Cao et al., 2023).
+But minimizing program size has obvious failure modes: code should also be understandable and extensible, which can be in tension with merely finding the shortest program.
 [^1] Perl Golf is a game where participants attempt to write the shortest Perl program accomplishing a given task.
 The resulting code is famously incomprehensible, even by the standards of Perl.
-To address these concerns, we follow prior work in program synthesis [16, 6, 17, 18] and define simplicity as the minimum description length (MDL), or negative log probability under a reference distribution.
-Formally, we are given a collection of code sources $\left\{\rho_{n}\right\}_{n=1}^{N}$, and output both a new library $\mathcal{L}$, as well as rewritten refactorings of the original code sources, $\left\{\rho^{\prime}_{n}\right\}_{n=1}^{N}$.
-We define the pass rate $\tau(\rho_{n})$ as the fraction of unit tests program $\rho_{n}$ passes.
-In practice we are concerned both with the case where we are refactoring several code sources ($N>1$) and also the case where there is only a single large code source we are refactoring ($N=1$).
-We optimize the following objective, which rewards refactorings that pass at least as many tests as the original program and minimize MDL:
-(,$\ell\left(\mathcal{L},\left\{\rho_{n}^{\prime}\right\}\right)=\begin{cases}-\log p_{\text{LM}}(\mathcal{L})+\sum_{n}-\log p_{\text{LM}}(\rho_{n}^{\prime}\mid\mathcal{L})&\forall\rho_{n},\tau(\rho_{n})\leq\tau(\rho_{n}^{\prime})\\ \infty&\text{otherwise}\end{cases}$ (1)
-where $p_{\text{LM}}(\rho_{n}^{\prime}|\mathcal{L})$ is the probability of the suffix $\rho_{n}^{\prime}$ given the prefix $\mathcal{L}$ under a language model, effectively concatenating the library and the program into one prompt, but only counting the perplexity of the later program tokens.
+Other work in program synthesis (Liang et al., 2010a; Solomonoff, 1964; Ellis et al., 2021) instead optimizes Minimum Description Length (MDL), or negative log probability under a reference distribution.
+In the software engineering community, other metrics such as cyclomatic complexity and maintainability index have been defined for similar purposes: These are more complex metrics that examine the syntax tree, call graph, and other statically-analyzable structures (McCabe, 1976).
+What metric should we use?
+We revisit this question in Section 6, where we empirically compare candidate metrics and human preferences before fixing our choice for the rest of the paper.
+For now assume a placeholder metric $M$ measuring refactoring quality; we seek to minimize $M$ while preserving correctness.
+Given a task comprising files $\left\{\rho_{n}\right\}_{n=1}^{N}$, we output both a new library $\mathcal{L}$, as well as rewritten refactorings of the original files, $\left\{\rho^{\prime}_{n}\right\}_{n=1}^{N}$.
+We define tests passed $\tau(\rho_{n})$ as the set of unit tests $\rho_{n}$ passes, and consider both refactoring several files ($N>1$) and also refactoring a single large file ($N=1$).
+We optimize the following objective, which prefers passing at least the same tests as the original programs and minimizing the chosen metric $M$:
+$\ell\left(\mathcal{L},\{\rho_{n}^{\prime}\}\right)=\begin{cases}M(\mathcal{L},\{\rho_{n}^{\prime}\})&\forall\rho_{n},\tau(\rho_{n})\subseteq\tau(\rho_{n}^{\prime})\\ \infty&\text{otherwise}\end{cases}$ (1)
+4 MiniCode—Library Design and Refactoring Benchmark
+MiniCode presents systems with a task comprising a set of files, then asks them to refactor the files into a unified library alongside refactorings of the original files.
+There are two key desiderata for benchmark tasks: They should have related files sharing latent abstractions, and should also be verifiable, to measure how well refactored files preserve functional correctness.
+We source a variety of problems (Table˜1).
+Table 1: MiniCode Statistics
+Domain Files Tasks Avg LoC Avg Tests / file
+Code Contests (Li et al., 2022b) 300 10 87 10
+Transformers (Wolf et al., 2020) 10 1 538 181
+Diffusers (von Platen et al., 2022) 11 2 685 75
+Logo (Wong et al., 2021) 300 1 10 1
+Date (Srivastava et al., 2023) 246 1 14 1
+CodeContests.
+Competition problems are crafted with specific variations of algorithmic approaches in mind, resulting in both shared latent concepts and the required test cases.
+As a result, competition coding is both verifiable, and ready to refactor.
+We therefore take solutions, prompts, and tests from CodeContests (Li et al., 2022a), a competition programming dataset.
+Huggingface Transformers Library.
+We test refactoring across implementations of large language and vision–language models from the Huggingface transformers repository (modelling_<name>.py files, e.g., Qwen2, LLaMA, DeepSeek-V3).
+Unlike competition coding, these sources are production-scale and Huggingface requires that all changes pass an extensive suite of integration tests before merging into the main branch.
+A refactoring is only deemed correct if it passes the unmodified Transformers test suite, making this a high-stakes setting that requires correctness and compatibility.
+Huggingface Diffusers Library.
+We test refactoring across implementations of diffusion models from the Huggingface diffusers repository (unet_<name>.py and scheduler_<name>.py files, e.g., Stable Diffusion UNet, DDPMScheduler), yielding two distinct tasks.
+Like Transformers, Diffusers requires that all changes pass a comprehensive suite of integration tests before merging into the main branch.
+Logo & Date.
+The library learning literature already has existing benchmarks: Typically they seek to learn a single library from a task comprising many sources, and then test that library on holdout program synthesis tasks.
+Logo and Date were used in the recent related work REGAL (Stengel-Eskin et al., 2024), which we incorporate wholesale to understand how our new method compares to state-of-the-art library learning.
+The associated programming problems were created by humans, but their solutions were generated by gpt-3.5-turbo.
 5 Librarian: Refactoring Code to Create Libraries
-This section details our method to compress collections of code sources into libraries, while migrating the code sources to use these shared building blocks.
-Figure 1 illustrates our method, Librarian.
-Librarian follows a simple sample-and-rerank framework to maximize our refactoring objective described in Section 3. It maintains and grows a library of useful functions as part of this objective.
-Concretely, our framework follows:
-,$\mathcal{L}^{\star},\left\{\rho^{\star}_{n}\right\}=\operatorname*{arg\,min}_{\mathcal{L},\left\{\rho^{\prime}_{n}\right\}\in\textsc{Sample}\left(\left\{\rho_{n}\right\}\right)}\ell\left(\mathcal{L},\left\{\rho^{\prime}_{n}\right\}\right).$ (2)
-4.1 Sample with clustering
-Meaningful abstractions exist primarily among programs that share some functionality or underlying structure.
-We perform clustering on the input programs to make groups of programs that likely share structures that can be abstracted into general library functions.
-Most modern language models cannot be prompted with the entire collection of input programs– even long context models cannot process the entirety of e.g the Linux kernel, and even if they could, it is not clear that such a strategy is the most efficient way of focusing the language model’s attention.
-We consider clustering algorithms for discovering small groups of related code; we call these tuples.
-This extends REGAL [3], which clusters programs solving similar problems by assuming each program is paired with a natural language description of the problem it solves, and clustering embeddings of those descriptions.
-But programs solving similar problems do not necessarily have similar structure.
-We therefore instead first summarize the code source itself by prompting a language model, then cluster based on the similarity of these new descriptions.
-Once identified, each tuple is used in two stages.
-The first is to retrieve relevant already-abstracted functions from the Librarian library: for a given tuple of programs, relevant functions are retrieved by prompting a language model with the entire existing existing library and the original input programs, then instructing it to identify which functions would be useful.
-The retrieved library functions and the original programs in the tuple are then provided as critical context to the language model to propose a sample budget of $K$ candidate refactorings.
-4.2 Rank with compression
-Once sampled, all $K$ candidate refactorings are passed through a sort-and-filter evaluation harness to select which one scores the highest on refactor quality and that maintains (or improves) test accuracy compared to the original.
-If no such candidate exists, the original code is preserved, maintaining existing functionality.
-New library functions in the selected refactor are saved into the Librarian library for potential use in downstream refactoring of other programs. We provide the full algorithm in Appendix A.
-5 MiniCode
-MiniCode evaluates a code agent’s capability to identify abstractions across implementations and design reusable libraries.
-In order to measure these capabilities, our benchmark presents agents with a collection of code sources, then asks agents to refactor the code sources into a unified library alongside refactorings of the original code sources.
-There are two key desiderata for collections of code sources: The collections must be compressible, in that there exists a latent shared library abstraction, and verifiable, so that we can measure how well refactored code sources preserve functional correctness.
-We source problems from two domains: Competition coding and synthesized repositories (Table 1).
-Agents are expected to interact with MiniCode via the terminal.
-We structure the benchmark as refactoring a multi-package Python repository, where each code source in a collection is a Python package in a subdirectory.
-This requires knowledge of basic bash commands for exploring repositories, editing code, and running tests, as well as how to manage complex, multi-package Python libraries.
-CodeContests
-Competition problems are crafted with specific variations of algorithmic approaches in mind, resulting in both shared latent concepts and required test cases.
-As a result, competition coding is naturally both compressible and verifiable.
-Each collection consists of multiple code sources, each containing a solution to a competition programming prompt, with associated tests for verification.
-We take solutions, prompts, and tests from CodeContests [19], a dataset consisting of competition coding problems.
-Each code source in the collection is structured as a subdirectory consisting of the task description in PROBLEM.md, the initial solution in main.py, and a script to run tests in run.sh.
-Agents are instructed to create a library.py file, which should is imported into each code source.
-Since CodeContests has no external dependencies on Python packages, this can be done without explicit structuring as a Python package.
-Repositories
-For the second domain of synthesized repositories, we propose a data-generating process that first produces project ideas, then generates variations of those project ideas tailored to specialized use-cases.
-This allows us to control the complexity and degree of overlap between each code source in a collection.
-Each code source in the repositories domain is comprised of a task description, source code, and test cases for functionality and correctness.
-MiniCode includes both small repositories (approximately 200 lines of source code each) and large repositories (approximately 6.5k lines of source code each), both of which represent realistic settings where different people with different needs use language models to help them write software for their particular use cases.
-The refactoring agent is tasked with extracting re-usable functions from across code repositories, and re-writing the original code source repositories to use them.
-We approximate the true distribution of code repositories, $p(\rho)$, with a generative process using latent textual library and repository descriptions.
-The entire space of code repositories is massive.
-To collect and group code repositories that share meaningful structure that can be abstracted into a useful code library, we (1) sample a textual library description; (2) sample use cases; and last (3) their programmatic implementations.
-This generative process naturally produces repository collections primed for refactoring.
-We sample from language models for each step of this process.
-Prompts are shared in Appendix E.
-Each collection consists of multiple synthetic repositories as code sources, where each code source is a subpackage.
-Collections are obtained by transforming the original repository code sources into a multi-package Python library.
-Each source’s source code is extracted into a subpackage directory, and its tests into a corresponding test subdirectory.
-Agents are instructed to write a shared library in a shared subpackage named common.
-This common shared library must be imported and used to refactor each of the original code sources.
-Table 1: MiniCode Statisics
-Domain Sources Collections Avg LoC Avg Tests Gen by
-Code Contests 300 30 87 10 Humans
-Small Repositories 262 22 209 12 o4-mini
-Large Repositories 20 10 6,433 101 Claude-Sonnet 3.7
-6 Experimental Setup
-Grouping Programs into Collections
-To facilitate parallel application of Librarian and manage the dataset scale, we assume that semantically distant code sources will have minimal overlap in their optimal library functions.
-Therefore, our overall approach partitions the dataset into disjoint collections through clustering.
-For CodeContests, these collections are constructed from an initial corpus of $\sim$9k problems with Python solutions: We first filter these code sources, removing those whose selected canonical solution is under 10 lines (minimal refactoring potential).
-For the remaining 4596 solutions we use a language model to generate textual descriptions of canonical solutions—emphasizing reusable components—which are embedded using OpenAI’s text-embedding-ada-002.
-Agglomerative Clustering [20] is subsequently applied to these embeddings to partition the code sources into a predefined number of initial clusters, in our case 120.
-To create uniformly sized experimental units, we subsample each such cluster to form collections of 30 code sources.
-This collection size was empirically chosen because it balanced between the runtime of Librarian without limiting compression.
-We select 10 collections that we then use to evaluate our methods.
-Code repositories are generated as disjoint collections through the generative process, which first samples project ideas then variations.
-For small repositories, we up sample $10-15$ code sources then form smaller collections of size $3$-$5$ code sources that exhibit the most intra-cluster similarity according to clustering using the same embedding and clustering technique.
-For big repositories, we sample two code sources for each of the $10$ collections.
-REGAL Baselines
-To evaluate the ability of our libraries to support reuse on new problems, we turn to the program synthesis tasks used in REGAL, where learned libraries are added to help the program synthesizer.
-We evaluate on the two domains published by the authors, Logo and Date.
-Because our clustering is inspired by REGAL but adds additional complexity, for fair comparison, we keep their setup the same and only augment the training using sample + MDL rerank procedure described in Section 4.1.
-Code Contests
-To evaluate Librarian on refactoring Code Contests we select 6 collections of 30 code sources (problems).
-In each collection we group the problems into tuples of size $3$.
-We set the sample budget to be $K=8$, since our ablations show that with larger $K$ we discover better libraries 3.
-We use the MDL objective for rankings.
-The model used for sampling is OpenAI’s o4-mini [21]. To obtain MDL scores we use Qwen 2.5 7B Instruct [22] as a balance between quality, speed, and cost.
-Code Agents
-To fairly evaluate performance on the task by state-of-the-art systems, we use coding agents that advertise long-context ability to reason about, write, and refactor code repositories.
-Specifically, we use Claude Code (Cl) [23] which uses the Claude 3.7 Sonnet model, and OpenAI’s Codex (Cx) [24] which uses o4-mini.
-All steps of the initial code sources for small code repositories are done by OpenAI Codex.
-We empirically observed that Claude Code has an affinity for generating large code repositories– the large code repositories are all generated by Claude Code.
-We test whether code agents can refactor collections of code sources autonomously, without human intervention.
-Refactoring repositories with code agents involves planning and iterative (re-)implementation and testing.
-Code agents are prompted to perform each of these steps, with feedback from the unit tests.
-Agents must run and repair unit tests autonomously.
-We run coding agents multiple times per collection, logging their progress in checklists stored in text files.
-Our naming convention for agents is “Planner-Executor”.
-For example, “Cl-Cx” uses the Claude Code agent for planning and OpenAI Codex to implement the plan.
-As before we use Qwen 2.5 7B Instruct [22] to obtain MDL scores.
+Librarian generates a new library from a set of files, while migrating the files to use that new library (Figure 1), following a sample-and-rerank framework: Prompting a backend LLM or agent to sample $K$ candidates, and picking the one minimizing the loss $\ell$.
+Naively,
+$\mathcal{L},\left\{\rho^{\prime}_{n}\right\}=\operatorname*{arg\,min}_{\mathcal{L},\left\{\rho^{\prime}_{n}\right\}\in\textsc{Sample}_{K}\left(\left\{\rho_{n}\right\}\right)}\ell\left(\mathcal{L},\left\{\rho^{\prime}_{n}\right\}\right)$ (2)
+for metric $M$ and sampling budget $K$.
+But this cannot work for large tasks with many programs, which would not fit into the context of most LLMs.
+Even long context models cannot process the entirety of e.g the Linux kernel, and even if they could, it is not clear that such a strategy is the most efficient way of focusing the language model’s attention.
+To address this, we wrap sample-and-rerank with a clustering algorithm that decomposes the task into manageable chunks, described next.
+Clustering.
+Meaningful abstractions arise when programs share underlying functionality or structure.
+To surface these, we cluster the task’s files into small groups that are likely to share reusable structure, and refactor each cluster separately from the rest.
+This decomposition shrinks the prompt size, and gives independent searches for the best per-cluster refactoring, which may be more tractable.
+Librarian’s clustering extends REGAL (Stengel-Eskin et al., 2024), which clusters programs by assuming each program is paired with a natural language description of the problem it solves, and clustering embeddings of those descriptions.
+Since similar problems need not imply similar solution code, we instead prompt a model to summarize each file and cluster by these summaries.
+Specifically, we define $\textsc{Cluster}_{S}\left(\left\{\rho_{n}\right\}\right)$ as performing agglomerative clustering (Ward Jr, 1963) on the task’s files $\left\{\rho_{n}\right\}$ to produce a set of set of files, each of which is a cluster of size $S$.
+We use text-embedding-ada-002 to embed descriptions of code sources for clustering.
+Combining clustering with sample-and-rerank.
+The simplest approach is to refactor each cluster independently and take the union (concatenation) of each cluster’s library:
+$\mathcal{L}^{\star}$ $=\bigcup_{c\in\textsc{Cluster}_{S}(\left\{\rho_{n}\right\})}\mathcal{L}_{c}$ (3)
+$\left\{\rho^{\prime}_{n}\right\}$ $=\bigcup_{c\in\textsc{Cluster}_{S}(\left\{\rho_{n}\right\})}\left\{\rho^{\prime}_{c,i}\right\}_{i=1}^{|c|}$ (4)
+$\mathcal{L}_{c},\left\{\rho^{\prime}_{c,i}\right\}_{i=1}^{|c|}$ $=\operatorname*{arg\,min}_{\mathcal{L},\left\{\rho^{\prime}_{c,i}\right\}_{i=1}^{|c|}\in\textsc{Sample}_{K}\left(c\right)}\ell\left(\mathcal{L},\left\{\rho^{\prime}_{c,i}\right\}_{i=1}^{|c|}\right)\text{, for each }c\in\textsc{Cluster}_{S}(\left\{\rho_{n}\right\})$ (5)
+The approach above ignores the fact that library abstractions discovered in one cluster might be useful in another cluster.
+A more sophisticated approach accumulates a library across clusters, and when refactoring a cluster, adds the accumulated library to the prompt.
+This lets abstractions discovered earlier carry forward across the collection.
+Appendix˜A describes this extension.
+6 What Makes a Good Refactoring?
+We compare different metrics $M$ measuring the quality of a refactoring:
+Tokens measures the total number of tokens in the refactored files and in the library.
+It minimizes program size, but not at the expense of creating a bloated library: Simply replacing every program with its own monolithic library function would not improve the tokens metric, because it measures library size as well.
+Concretely, $M_{\text{tokens}}(\mathcal{L},\{\rho^{\prime}_{n}\})=\textsc{Tokens}(\mathcal{L})+\sum_{n}\textsc{Tokens}(\rho^{\prime}_{n})$.
+Minimum Description Length (MDL) evaluates the negative log probability under a reference distribution, taking into account both the library and refactored sources.
+Concretely, $M_{\text{MDL}}(\mathcal{L},\{\rho^{\prime}_{n}\})=-\log p_{\text{LM}}(\mathcal{L})+\sum_{n}-\log p_{\text{LM}}(\rho^{\prime}_{n}\mid\mathcal{L})$, where $p_{LM}(\rho_{n}^{\prime}|\mathcal{L})$ is concatenating the library and the program into one prompt, but only counting the perplexity of the later program tokens.
+This has a Bayesian justification: The MDL library is the maximum aposteriori estimate of $\mathcal{L}$ given conditionally-independent programs.
+We use Qwen-2.5-3B as our reference language model, as it is modern, open, and has publicly-available endpoints for querying logits, which is required for scoring refactorings.
+To confirm that our MDL optimization results are not model specific, we computed MDL values for 15 Code Contests clusters, each of 50 valid refactoring candidates, using Qwen and Llama-3.2-3B models.
+We found a 94% agreement in the minimum MDL candidates from both models.
+Cyclomatic Complexity (CC) is a longstanding metric from the software engineering community which measures the number of linearly independent paths through a program’s control flow graph (McCabe, 1976).
+Smaller programs often have lower cyclomatic complexity.
+It is equivalent to defining $M_{\text{CC}}(\mathcal{L},\{\rho^{\prime}_{n}\})=\textsc{cc}(\mathcal{L})+\sum_{n}\textsc{cc}(\rho^{\prime}_{n})$ and $\textsc{cc}(\rho)=E(\rho)-N(\rho)+2P(\rho)$, where $E$, NN, and $P$ measure the number of control flow edges, nodes, and connected components, respectively.
+Maintainability Index (MI) is a modern software engineering metric combining several other metrics, including lines-of-code, cyclomatic complexity, and Halstead volume into a single score.
+Higher MI values are intended to indicate easier-to-maintain code, so we define $M_{\text{MI}}(\mathcal{L},\{\rho^{\prime}_{n}\})=-\textsc{MI}(\mathcal{L})+\sum_{n}-\textsc{MI}(\rho^{\prime}_{n})$.
+6.1 Asymptotic behavior of metrics in large-sample regime
+Are these metrics equally effective at encouraging modular and reusable libraries?
+To answer this question, we run Librarian on 15 CodeContests (each of three files) using MDL, tokens, maintainability index, and cyclomatic complexity, while varying the inference-time sample budget $K$ (Figure˜2,7).
+We use Best@k estimator for the expected value of metrics for all $k\leq K$ (we describe the estimator and prove its correctness in Appendix F).
+Tokens and MDL separate cleanly from classic software engineering metrics: Optimizing tokens/MDL, both of which essentially compress the original programs, does not yield steady improvements in MI/CC, and vice-versa.
+To understand whether these libraries expose shared abstractions, we examine the average number of times that each library routine is used, and the average number of library invocations per library function.
+This teases apart tokens and MDL: Optimizing MDL yields more more reusable libraries (used about $8\times$ per task), with each function called more often (called about $2.2\times$ per function)—exceeding the other metrics we consider.
+See Appendix Figure˜7 for full results, including Cyclomatic Complexity.
+Figure 2: (A) Asymptotic behavior of metrics for scoring libraries and refactorings (columns) varying refactoring budget (horizontal axes).
+(B) Comparing metrics via proxies of downstream library quality (total library usage and average calls per library function), for which MDL>>Tokens>>MI.
+All results are estimated using Best@k.
+See also Appendix Figure˜7.
+Figure 3: Best@K MDL ratio. Increasing sample budget improves MDL on Transformers.
+Studying these metrics at large $k$ allows understanding their inference-time scaling behavior.
+While the underlying metric itself improves (with diminishing returns), this was not to be taken for granted: The backend language model must produce sufficient diversity to steadily improve these metrics.
+Prior state-of-the-art, such as Stengel-Eskin et al.
+(2024), instead take a single sample, but the results here suggest benefit from further test-time search, and indeed, real-world repos benefit from steady improvement with increased samples (Figure 3).
+But our proxies of library utility plateau much earlier, around $k=20$ samples, suggesting large $k$ is unnecessary in practice: Effective library building benefits from test-time compute, but does not demand an exorbitant amount of it.
+6.2 Human evaluation of refactoring metrics
+Figure 4: Human evaluation of different refactoring objectives. Judges compare pairs of refactorings that both pass all test cases. MDL aligns best with human preferences.
+We perform a human study to corroborate the findings of Section˜6.1 using the exact same CodeContests clusters (Appendix˜E).
+The human study compares tokens, MDL, and Maintainability Index by (1) refactoring clusters into libraries, (2) presenting human participants with the original sources and their refactorings under pairs of metrics, and (3) eliciting pairwise preferences from human participants.
+Humans prefer MDL-minimizing libraries (Figure˜4), and although the preference is only statistically significant for MDL vs.
+MI, the data suggest a rank-order preference of MDL>Tokens>MI; Tokens vs.
+MI was slightly below the threshold of significance.
+We ran $N=14$ participants (eliciting 129 judgments), and although more participants would make more of these comparisons statistically significant, already we see a general preference for compression-based metrics (MDL and Tokens) with only MDL crossing the threshold of statistical significance.
+Indeed, minimizing MDL is not quite the same as minimizing tokens, as Figure˜5 illustrates: We believe basically every human coder would prefer the MDL-minimizing program in this example.
+We therefore adopt $M_{MDL}$ as the primary objective in the remainder of this paper: In addition to support from this human study, (1) Bayesian arguments support MDL; (2) corner cases in the style of ‘Perl golf’ provide existence proofs of the liability of merely minimizing tokens; and (3) reasonable proxies for library reuse favor MDL (Section˜6.1).
+Figure 5: Example where tokens and MDL diverge: Obfuscating the original library definitions (left) by shortening variable names (right) reduces tokens but increases MDL.
+7 What we learn from running Librarian on MiniCode
+We empirically study Librarian on MiniCode with the goal of understanding (1) the degree to which library abstractions are reused across programs, (2) how our method compares to state-of-the-art library learning on existing datasets, and (3) whether Librarian holds value for real-world repos.
+Librarian discovers reusable functions for competition programming–but some functions are only called once.
+We test on CodeContests with a cluster size of $S=3$ and a sample budget of $K=8$ draws from o4-mini, as reasoning models perform well on competition programming.
+[^2] Code agents such as Codex, Claude Code, and others underperformed o4-mini (Appendix I) Table˜2 shows that the resulting refactors and libraries approximately halve the MDL, which incidentally reduces program size as well (44% relative reduction in token count).
+Pass rate modestly improves as an incidental consequence of sampling and filtering with test cases.
+Libraries average 10 functions, each heavily reused: Averaging 5 uses per function within tasks comprising only 10 programs.
+But almost 40% of library functions are only used once.
+Why is that?
+A signature of the MDL objective is a preference for whatever a language model assigns high apriori probability to.
+Although a single-use function does not reduce line count or tokens —the function could simply be inlined—it improves MDL if it yields a more natural decomposition of the target programs.
+Indeed, human-written libraries sometimes include functions that are seldom used, provided they serve as a conceptually modular abstraction.
+We therefore see single-use functions as a feature, not a bug.
+See Appendix J for an example refactoring candidate on CodeContests.
+Are these libraries useful for solving new, unseen programming problems?
+Library learning has long sought to learn libraries from training programs which then help solve new unseen program synthesis tasks.
+The Logo and Date datasets fit within this paradigm.
+Recently REGAL improved the state-of-the-art on these library learning datasets.
+Because our clustering is heavily inspired by REGAL, for fair comparison, we keep exactly their clustering setup but add MDL-based reranking using $K=5$ samples.
+Despite the simplicity of these datasets, we find value in our more complicated method.
+Table˜3 shows that sampling and reranking by MDL yields up to a 41.8% relative improvement in solve rate on unseen programming problems, and that even when the gains are more modest, we still improve upon the state-of-the-art.
+But these are relatively simple problems solvable with about ten lines of code—does this work in the real world?
+Table 2: Results for Librarian on 10 Code Contests tasks ($K=8,S=3$)
 Metric Value
@@
 Avg Calls per Function 5.17 ±1.08
-Percentage Single Use Functions 38.03% ±4.88
-Figure 2: Refactoring results for Librarian (w/ $K=8$) averaged over 10 Code Contests collections.
-7 Results
-In this section, we present the compression and correctness results with Librarian and agent baselines on MiniCode.
-MiniCode-CodeContests On CodeContests, Librarian achieves a high final pass rate of 90.67% and significantly improves correctness, with pass rates increasing by 6.33% compared to the original code sources (Table 2).
-The method yields substantial compression: the refactored code, including the new library, shows an MDL ratio of 0.53 (a 47% reduction in MDL relative to the original).
-On average, Librarian generates libraries containing approximately 11 functions.
-These functions demonstrate good reuse, being called by around 5.2 programs on average, although 38.03% of them are used only once within their specific collection context.
-Code agents fail to achieve both high correctness and compression on MiniCode-CodeContests.
-Across collections, the Codex agent achieves an average MDL ratio of 0.83, but a pass rate of 74.16%, much lower than Librarian’s rate of 90.67%.
-Similarly, the Claude agent reaches a higher pass rate of 82.50%, which is still lower than Librarian, but an MDL ratio of 1.04 which is more complex than the original collection.
-MiniCode-Repositories Coding agents results on both small and large repositories are given in Table 2.
-During experiments, we found that the Claude Code agent provided superior plans for refactoring, while both the Claude Code and Codex agents performed satisfactorily at implementing a plan.
-Despite using code agents with a human in the loop, existing code agents fail to produce effective refactorings.
-Often, the resulting refactored codebase bloats in size and complexity, particularly when using Claude Code for both planning and implementation.
-Given the poor performance of code agents on these repository-scale refactorings, we do not run the full sample-and-rerank pipeline.
-Supporting qualitative analysis in Section 8.4 discusses common failure modes of code agents on this task in repository-size refactorings.
-Collection Agent MDLPass%
-datapipe orig 1.0 1.0
-Cl-Cl 9.1 1.0
-Cl-Cx 2.2 1.0
-state_machine orig 1.0 1.0
-Cl-Cl 6.3 0.9
-Cl-Cx 1.9 1.0
-config_schema orig 1.0 1.0
-Cl-Cl 2.5 0
-Cl-Cx 2.1 1.0
-cli_tools orig 1.0 1.0
-Cl-Cl 4.0 1.0
-Cl-Cx 3.1 0
-cli_form orig 1.0 1.0
-Cl-Cl 3.8 1.0
-Cl-Cx 2.9 1.0
-Collection MDLPass%
-filesys_analyzer orig 1.0 1.0
-Cl-Cl 1.6 1.0
-query_language orig 1.0 1.0
-Cl-Cl 1.4 1.0
-knowledge_store orig 1.0 1.0
-Cl-Cl 1.4 1.0
-vm_emulator orig 1.0 1.0
-Cl-Cl 1.6 1.0
-finance_tracker orig 1.0 1.0
-Cl-Cl 1.8 0.6
-in_memory_db orig 1.0 1.0
-Cl-Cl 1.5 1.0
-text_editor orig 1.0 1.0
-Cl-Cl 1.2 1.0
-Table 2: The correctness (Pass %) and compression (MDL ratio %) of the original and refactored code sources for the agent baselines on a subset of the small and large repository collections in MiniCode.
-Libraries with syntax or runtime errors receive a pass rate of 0.
-8 Analysis
-We further analyze parts of Librarian, consider alternative objective functions, analyze the use of clustering while constructing MiniCode-CodeContests, and common failure modes of code agents.
-8.1 Can these libraries be reused to solve downstream programming problems?
-To understand the practical value of our library extraction work for program synthesis, we augment the existing library learning algorithm, REGAL with sampling and MDL-reranking.
-In REGAL, the learned library is used to solve downstream holdout program synthesis problems, and a simpler clustering algorithm is used.
-For fair comparison with REGAL we hold the clustering algorithm constant (using the simpler REGAL approach), but sample $K=5$ refactorings/libraries and rerank by MDL.
-The resulting system outperforms REGAL on holdout programming problems (Table 3).
-These problems however are highly simple, which is representative of classic library learning work: Either drawing simple geometric designs (dataset Logo) or manipulating textual representations of dates (dataset Date).
-This helped motivate us to consider refactoring whole repositories.
-Table 3: Solving downstream program synthesis tasks using learned libraries
+% Single Use Functions 38.03% ±4.88
+Table 3: Solving holdout test program synthesis tasks using learned libraries.
 Dataset Model Pass Rate
@@
 Librarian (3.5-turbo) 94.7% ±0.7
-8.2 What objective function should a library learner use?
-We run our method on CodeContests using MDL, number of tokens, and cyclomatic complexity [25] [^2] Cyclomatic complexity (CC) is a classic metric from the software engineering community that analyses the number of paths through a program’s control flow, and unlike MDL is only loosely coupled with program size.
-as objective functions on 6 collections (Figure 3) While minimizing MDL also minimizes the other two objectives, the converse is not true.
-This suggests that MDL is a pareto-optimal loss among the three objectives in this experiment.
-To confirm that the library does indeed expose shared abstractions, we calculate the average number of times that each library routine is used.
-Scaling the inference budget to $K=8$ discovers better libraries, reusing each library function on average about 5 times.
-In Appendix A.1, we find that MDL is also the objective that best correlates with human preferences in a small-scale human study.
-Figure 3: Comparing 3 different objective functions for refactoring (different columns) according to different downstream success metrics (different rows), as a function of refactoring budget (horizontal axes).The values are averaged over 6 collections of CodeContests problems.
-Row 1: Optimizing perplexity also incidentally optimizes cyclomatic complexity and token count, but that the converse is not true.
-Row 2: refactored programs pass more test cases, even more than the original code itself.
-Row 3: increasing the refactoring budget results in more reusable library subroutines (such subroutines are called more times on average).
-Filtered/Raw: Using/Not Using tests to filter samples.
-8.3 Clustering Analysis: CodeContests
-We analyze the coherence of the clusters underlying collections in MiniCode-CodeContests.
-In particular, we compare clustering based on o4-mini generated code source descriptions against task descriptions.
-Since task descriptions in competition coding problems are designed to hide the algorithmic approach needed to solve problem, we expect that clusters based on code source descriptions are more coherent.
-We use two measures to evaluate the thematic coherence of collections: Good collections should group code sources with a (1) concentrated and (2) identifiable set of shared conceptual tags, which for CodeContests are provided as ground truth (trees, graphs, etc.).
-Normalized Tag Instance Entropy
-This measures the concentration of tag instances within a collection, and is given by the entropy of the tag distribution for a given collection, normalized by the number of distinct tags in that collection.
-A lower normalized tag instance entropy (closer to 0) indicates higher thematic purity, meaning a small number of tag types are most prevalent.
-Herfindahl-Hirschman Index (HHI) for Problem Presence
-This measures tag concentration across distinct code sources in a collection. A higher HHI signifies that the problems are collectively characterized by a smaller, more focused set of tags.
-We provide the full formal definitions of both measures in Appendix B.
-Figure 4 shows our clustering approach yields more thematically coherent clusters, evidenced by achieving lower entropy and higher HHI values across the entire tested range of $N$.
-Figure 4: Clustering analysis of 4,596 Code Contest problems, comparing the thematic coherence of clusters formed using our proposed method versus REGAL-style clustering.
-8.4 Common Failure Modes of Coding Agents
-Failing to follow the plan.
-Refactoring large repositories requires editing code over a long time horizon.
-We instruct agents to plan hierarchically by first recording their plan in natural language, then implementing that plan.
-However, agents often ignore the plan, particularly with Codex as the implementation agent, potentially refusing to edit the code at all.
-We hypothesize that this behaviour is caused by the refactoring task itself – the original code source passes all tests, and agents are trained to prioritize correctness over compression.
-Failing to use the library.
-When the implementation agent does follow the plan, it often fails to do so effectively.
-Claude Code, for example, will write a library, import it to the proper place, then rewrite the library rather than use it.
-We provide further examples in Appendix G.
+Real-World Refactoring.
+The HuggingFace Transformers library is used by nearly 400k GitHub projects.
+We deploy Librarian to 10 source files, using Claude Code to sample $K=15$ refactorings per cluster of size $S=5$, believing an agent such as Claude Code would excel at repo-level edits.
+Librarian distilled repeated abstractions such as MLPs, Attention, Decoder classes, RoPE helper functions, etc., lowering MDL to 67.2% of its original value while still passing all integration tests.
+The top-3 refactorings based on MDL have an average of $18\pm 4.4$ abstractions (functions, classes) in the library, each of which is called on average $4.59\pm 0.39$ times in the refactored models.
+For Diffusers, scheduler clusters yielded top-3 MDL refactorings with an average of $12.3\pm 1.6$ functions and $3.0\pm 0.4$ calls per function, while UNet refactorings produced richer abstractions with an average of $17.0\pm 5.6$ functions/classes and $3.43\pm 0.67$ calls each.
+Refactoring at scale proved expensive: Each refactoring took approximately 30 minutes to generate and test.
+But this is a one-off cost, and in our view, the refactored Transformers and Diffusers sources are much cleaner, and the new library is transparently reusable (Figure˜6).
+To the best of our knowledge, this is the first time any library learning algorithm has been successfully applied to real-world software projects.
+Figure 6: Representative result for refactoring HuggingFace Transformers using Librarian
+Learned libraries from these real-world codebases are useful for unseen downstream refactoring tasks.
+When a library learned on one cluster of Transformer files (5 models) is applied to refactor a second cluster, Librarian reduces the unseen cluster’s MDL to 73% of the its original value, with an average of 3.0 calls per library function.
+This demonstrates that Librarian learned libraries that can be repurposed to more compactly rewrite unseen real-world code sources.
+Human performance.
+HuggingFace is concurrently refactoring the Transformers library.
+On the same 10 source files used above, their refactoring achieves an MDL compression ratio of 66.5%.
+To estimate an independent ceiling, we the authors refactor the task ourselves and obtain 62%.
+This suggests room for a little improvement beyond Librarian’s 67.2%, but not by much (from inspection, specifically with better implementations of mixture-of-experts and vision models).
 8 Conclusion
-We introduce a new benchmark MiniCode and method Librarian for compressing code sources through reusable abstractions.
-We highlight the challenges that modern models face when producing modular and maintainable code, then present an effective method for using LLMs to do this task in small-scale programs.
-By framing refactoring as both a design and compression task, our work opens new directions for building more general and scalable code understanding and generation systems.
-In particular, the structure of MiniCode lends itself well to a reinforcement learning-based task, which synthesizes collections of repositories to refactor then computes rewards based on compression.
-Limitations Limitations of this work include the evaluation of synthetic repositories, the poor performance of code agents on refactoring, and the fact that compression may not be correlated with reuse.
-While our experiments on downstream programming problems partly address the question of reuse, investigating how to measure and encourage reuse for large-scale multi-repo library creation remains an open problem.
-Ethics Statement
-As with any coding benchmark, our work has the potential to improve automated coding.
-However, this has the potential to produce malicious code if left to run unchecked.
-Care should be taken to monitor inputs and outputs to the system.
+We introduce a new benchmark MiniCode and method Librarian for compressing files through reusable abstractions.
+We highlight the challenges of producing modular and maintainable libraries, then present an effective method for using LLMs to do this task.
+By framing refactoring as an optimization problem, our work opens new directions for building more general and scalable code understanding and generation systems.
+In particular, the structure of MiniCode lends itself well to reinforcement learning, where training would entail synthesizing collections of repositories to refactor then computing rewards based on MDL or other metrics.
+Limitations.
+We evaluate on synthetic toy problems (Logo and Date), and on competition programming problems, neither of which are naturalistic, although this is partly counterbalanced by our study of real-world refactoring.
+Compression, whether measured by tokens or MDL, may not always correlate with reuse, a limitation we sought to address through our experiments on downstream programming problems, and on holdout Transformers files: But investigating reuse on unseen app-building problems for real-world repo-level refactors remains open.
 Acknowledgements
-Appendix A Algorithm
+KE and ZC were supported by NSF grant #2310350.
+Appendix A Incremental Version of Librarian
+We support sharing library components across clusters through the below incremental version of our algorithm.
+It processes clusters sequentially, rather than in parallel, and its output can depend on the ordering imposed upon the clusters.
+$\mathcal{L}_{0}=\varnothing$
+$\{\rho_{n}^{\prime}\}=\bigcup_{c\in\text{CLUSTER}(\{\rho_{n}\})}\{\rho_{c,i}^{\prime}\}_{i=1}^{|c|}$
+$(\Delta\mathcal{L}_{c},\{\rho_{c,i}^{\prime}\}_{i=1}^{|c|})=\arg\min_{(\Delta\mathcal{L},\{\rho_{i}^{\prime}\})\in\text{SAMPLE}_{k}(c;\,\mathcal{L}_{t-1})}\ \ell(\mathcal{L}_{t-1}\cup\Delta\mathcal{L},\{\rho_{i}^{\prime}\}_{i=1}^{|c|})\quad\forall c\in\text{CLUSTER}(\{\rho_{n}\})$
+$\mathcal{L}_{t}=\mathcal{L}_{t-1}\cup\Delta\mathcal{L}_{t},\qquad\mathcal{L}^{\star}=\mathcal{L}_{|\text{CLUSTER}(\{\rho_{n}\})|}$
+Appendix B Algorithm
 Algorithm 1 Refactoring Specialized Programs into a Joint Library
@@
 3:Joint library $\mathcal{L}_{final}$ and set of refactored programs $P_{final}$
-4:$\mathcal{C}\leftarrow\text{Cluster}(P_{initial})$
-5:$\mathcal{L}_{final}\leftarrow\emptyset$, $P_{final}\leftarrow\emptyset$
+[…]
+[…]
 6:for all cluster $c\in\mathcal{C}$ do $\triangleright$ Each cluster independently
-7:     $T_{C}\leftarrow\text{GroupIntoTuples}(c)$triangleright Get tuples for each cluster
+7:  $T_{C}\leftarrow\text{GroupIntoTuples}(c)$ $\triangleright$ Get tuples for each cluster
 8:  for all tuples $\tau\in T_{C}$ do
@@
 11:   for $i=1\text{ to }K$ do $\triangleright$ Sample $k$ times
-12:              ($(\{f_{new,i}\},\{\rho^{\prime}_{i}\}\})\leftarrow\textsc{Sample}\left(f_{retrieved},\tau\right)$
+12:     $(\{f_{new,i}\},\{\rho^{\prime}_{i}\}\})\leftarrow\textsc{Sample}\left(f_{retrieved},\tau\right)$
 13:     $S\leftarrow S\cup\{(\{f_{new,i}\},\{\rho^{\prime}_{i}\})\}$
 14:   end for
-15:         $(f_{best},\{\rho^{\prime}_{best}\})\leftarrow\text{RerankAndSelectBest}(S,\ell\left(\cdot\right))$triangleright Rerank using objective
+15:   $(f_{best},\{\rho^{\prime}_{best}\})\leftarrow\text{RerankAndSelectBest}(S,\ell\left(\cdot\right))$ $\triangleright$ Rerank using objective
 16:   $\mathcal{L}_{final}\cup\{f_{best}\}$
@@
 20:return $\mathcal{L}_{final},P_{final}$
-A.1 Objective Functions
-Human evaluation best aligns with the MDL objective function.
-We collect 19 CodeContests problems and a pair of refactorings for each, and have 4 human judges pick which they prefer.
-Each refactoring comprises three rewritten programs, plus the library generated by Librarian.
-Each pair of refactorings passed all testcases, and were chosen to include both the highest and lowest perplexity refactorings.
-Human judges are authors on this paper, but were ‘blinded’ to not know relative objective function scores for the candidate refactorings.
-An MDL objective aligns best with human preferences (Figure 5), although number of tokens is also a reasonable proxy for human judgment.
-That token count and MDL yield different statistics furthermore implies that the lowest MDL program is not always the shortest program.
-Figure 5: Human evaluation of different refactoring objectives. Humans and models give preferences to pairs of refactorings/libraries.
+Appendix C Experimental Setup
+Grouping Programs into Collections
+To facilitate parallel application of Librarian and manage the dataset scale, we assume that semantically distant files will have minimal overlap in their optimal library functions.
+Therefore, our overall approach partitions the dataset into disjoint collections through clustering.
+For CodeContests, these collections are constructed from an initial corpus of $\sim$9k problems with Python solutions: We first filter these files, removing those whose selected canonical solution is under 10 lines (minimal refactoring potential).
+For the remaining 4596 solutions we use a language model to generate textual descriptions of canonical solutions—emphasizing reusable components—which are embedded using OpenAI’s text-embedding-ada-002.
+Agglomerative Clustering (Ward Jr, 1963) is subsequently applied to these embeddings to partition the files into a predefined number of initial clusters, in our case 120.
+To create uniformly sized experimental units, we subsample each such cluster to form collections of 30 files.
+This collection size was empirically chosen because it balanced between the runtime of Librarian without limiting compression.
+We select 10 collections that we then use to evaluate our methods.
+For Transformers, since the number of models is on the lower end, we manually chose a set of popular LLM / VLM models and passed them to the agent in collections of 5 code sources.
+REGAL Baselines.
+To evaluate the ability of our libraries to support reuse on new problems, we turn to the program synthesis tasks used in REGAL, where learned libraries are added to help the program synthesizer.
+We evaluate on the two domains published by the authors, Logo and Date.
+Because our clustering is inspired by REGAL but adds additional complexity, for fair comparison, we keep their setup the same and only augment the training using sample + MDL rerank procedure described in Section˜5.
+Code Contests.
+To evaluate Librarian on refactoring Code Contests we select 6 collections of 30 files (problems).
+In each collection we group the problems into tuples of size $3$.
+We set the sample budget to be $K=8$, since our ablations show that with larger $K$ we discover better libraries 2.
+We use the MDL objective for rankings.
+The model used for sampling is OpenAI’s o4-mini (OpenAI, 2024).
+To obtain MDL scores we use Qwen 2.5 7B Instruct (Qwen et al., 2025) as a balance between quality, speed, and cost.
+Code Agents on Transformers and Diffusers Repositories.
+To fairly evaluate performance on the task by state-of-the-art systems, we use coding agents that advertise long-context ability to reason about, write, and refactor code repositories.
+Specifically, we use Claude Code (Cl) (Anthropic, 2025) which uses the Opus 4.1.
+We test whether code agents can refactor collections of code sources autonomously, without human intervention.
+Refactoring repositories with code agents involves planning and iterative (re-)implementation and testing.
+Code agents are prompted to perform each of these steps, with feedback from the unit tests.
+Agents must run and repair unit tests autonomously.
+We run coding agents multiple times per task, logging their progress in checklists stored in text files.
 The instruction provided to human evaluators is as follows:
@@
 Listing 1: Human Evaluation Instruction
-Appendix B Collection Coherence Measures
-We provide the full definitions of the collection coherence measures.
+Appendix D Further Results on Refactoring Metrics
+Here we include the full graph of asymptotic behavior of all four scoring metrics (MDL, tokens, MI and CC), reporting the raw library metrics and the metric ratios.
+We can see that refactored programs have higher CC than the baseline and that optimizing CC as an objective does not decrese the other metrics.
+MDL performs best on total raw library usage as well as on average times a library function is used in the refactorings.
+MI ends up optimizing the number of library functions the best, but their reusability is below the average for the refactorings.
+Optimizing for tokens produces smaller libraries with less usage per function compared to optimizing MDL.
+Figure 7: Asymptotic behavior of metrics for scoring libraries and refactorings (columns) varying refactoring budget (horizontal axes).
+Appendix E Human Study Details
+We ran a user study with 14 participants where each participant had to judge 10 refactoring pairs.
+Each pair was comparing two out of three metrics, MDL, tokens, and MI.
+We showed the participants the original programs, as well as both of the refactoring versions (including the programs and the learned libraries).
+The participants were able to choose either version or say that the two refactorings are almost the same.
+To quantify pairwise preferences between refactoring metrics, we employed the Bradley-Terry model, a standard framework for analyzing paired comparison data.
+We fit the model using maximum likelihood estimation with mutual information (MI) as the reference category $(\pi_{M}I=1)$.
+To address potential noise in human judgments, we applied consensus-based filtering with a 75% threshold, retaining all responses for comparisons with low consensus (indicating genuine ambiguity) while excluding minority responses on high-consensus comparisons where the majority preference likely indicates the correct judgment.
+This conservative approach preserved 92.4% of responses while strengthening the statistical evidence for metric preferences, with MDL significantly outperforming MI ($p=0.67$, $95\%$ Confidence Interval: $[0.56,0.78]$).
+Even without the filtering MDL preference over MI was statistically significant.
+Appendix F Best@k Compression is a U-Statistic
+We wish to estimate the expected compression ratio achieved by our sample + rerank method, which samples $k$ candidate refactorings, discards any that do not pass the tests, and selects the one with the lowest score (total log-prob).
+Background on U-Statistics. Let $Z_{1},\dots,Z_{n}\overset{\text{i.i.d.}}{\sim}F$. For a symmetric function $h:\mathcal{Z}^{k}\to\mathbb{R}$, the U-statistic of order $k$ is defined as
+$U_{n}\;=\;\binom{n}{k}^{-1}\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}h(Z_{i_{1}},\dots,Z_{i_{k}}).$ (6)
+By construction,
+$\mathbb{E}[U_{n}]\;=\;\mathbb{E}[h(Z_{1},\dots,Z_{k})],$ (7)
+so $U_{n}$ is an unbiased estimator of the population quantity $\theta=\mathbb{E}[h(Z_{1},\dots,Z_{k})]$.
+Application to Best@k Compression. Let each valid refactoring be a pair $Z=(S,C)$, where $S$ is the score and $C$ is the compression ratio. Define the symmetric function
+$h_{k}(z_{1},\dots,z_{k})\;=\;C_{j^{*}},\qquad j^{*}=\arg\min_{1\leq j\leq k}S_{j},$ (8)
+the compression ratio of the lowest-score refactoring among $k$ draws. The population target is then
+$\theta_{k}=\mathbb{E}[h_{k}(Z_{1},\dots,Z_{k})].$ (9)
+Given $n$ valid samples, our estimator is
+$\widehat{\theta}_{k}\;=\;\binom{n}{k}^{-1}\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}h_{k}(Z_{i_{1}},\dots,Z_{i_{k}}).$ (10)
+Proposition. $\widehat{\theta}_{k}$ is a U-statistic of order $k$ with function $h_{k}$, and hence an unbiased estimator of $\theta_{k}$.
+Proof.
+(1) Symmetry of the function $h_{k}$.
+$h_{k}$ selects the compression associated with the lowest score among its $k$ arguments.
+Permuting the inputs does not affect this outcome (ties can be resolved with a fixed, permutation-invariant rule).
+Thus $h_{k}$ is symmetric.
+(2) U-statistic form. By definition, a U-statistic of order $k$ with kernel $h_{k}$ is
+$U_{n}=\binom{n}{k}^{-1}\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}h_{k}(Z_{i_{1}},\dots,Z_{i_{k}}),$
+which matches $\widehat{\theta}_{k}$ exactly.
+Therefore, $\widehat{\theta}_{k}$ is a U-statistic of order $k$. By the unbiasedness property of U-statistics,
+$\mathbb{E}[\widehat{\theta}_{k}]=\theta_{k}.$
+$\square$
+Thus, our reported best@k compression curves provide unbiased estimates of the expected performance of the sample + rerank method.
+Appendix G Clustering Analysis: CodeContests
+We analyze the coherence of the clusters underlying collections in MiniCode-CodeContests.
+In particular, we compare clustering based on o4-mini generated file descriptions against task descriptions.
+Since task descriptions in competition coding problems are designed to hide the algorithmic approach needed to solve problem, we expect that clusters based on file descriptions are more coherent.
+We use Normalized Tag Instance Entropy and Herfindahl-Hirschman Index to evaluate clusterings.
+Figure˜8 shows our clustering approach yields more thematically coherent clusters, evidenced by achieving lower entropy and higher HHI values across the entire tested range of $N$.
+We provide definitions of our measures below.
+Figure 8: Clustering analysis of 4,596 Code Contest problems, comparing the thematic coherence of clusters formed using our proposed method versus REGAL-style clustering.
+G.1 Collection Coherence Measures
+We use two measures to evaluate the thematic coherence of collections: Good collections should group files with a (1) concentrated and (2) identifiable set of shared conceptual tags, which for CodeContests are provided as ground truth (trees, graphs, etc.).
+We provide the full definitions of the collection coherence measures below.
 Normalized Tag Instance Entropy:
@@
 where $\text{Tags}(C)$ represents the set of unique tags present in cluster $C$.
-Appendix C Benchmark Comparison
+Appendix H Benchmark Comparison
 We compare our benchmark, MiniCode, to similar benchmarks in Table 4.
@@
 For example, optimizing for program correctness alone does not require exploring a large solutions space, whereas optimizing a program for speed would.
-In the case of compressing large code sources, we must explore the large space of shared abstractions afforded by libraries in order to maximize compression.
+In the case of compressing large files, we must explore the large space of shared abstractions afforded by libraries in order to maximize compression.
 Table 4: Comparison of Code Benchmarks
 Benchmark Creativity/Design Scale
-SWE-bench [9] Low Repository
-Commit-0 [10] Medium Repository
-RefactorBench [11] Low File
-ECCO [13] High Function
-KernelBench [14] High Function
-MiniCode(Ours) High Multi-repository
-Appendix D Full MiniCode Repository results
-We present the full repository-level results of MiniCode-repository in Tables 5 and 6.
-Collection Agent LLoC CC MDL ratio Pass rate
-datapipe original 474 116 1.0 1.0
-Cl-Cl 1084 341 9.1 1.0
-Cl-Cx 645 174 2.2 1.0
-state_machine original 619 222 1.0 1.0
-Cl-Cl 2271 735 6.3 0.9
-Cl-Cx 686 227 1.9 1.0
-config_schema original 949 284 1.0 1.0
-Cl-Cl 922 339 2.5 failed
-Cl-Cx 626 207 2.1 1.0
-cli_tools original 875 300 1.0 1.0
-Cl-Cl 2925 909 4.0 1.0
-Cl-Cx 5848 2378 3.1 failed
-cli_form original 352 174 1.0 1.0
-Cl-Cl 855 342 3.8 1.0
-Cl-Cx 1366 548 2.9 1.0
-Table 5: Full results on MiniCode-repositories small.
-Collection LLoC CC MDL ratio Pass rate
-file_system_analyzer original 3911 1338 1.0 1.0
-refactored 6495 2218 1.6 1.0
-query_language_interpreter original 5181 2105 1.0 1.0
-refactored 7193 2966 1.4 1.0
-personal_knowledge_management original 5341 1832 1.0 1.0
-refactored 6357 2364 1.4 1.0
-virtual_machine_emulator original 6841 2283 1.0 1.0
-refactored 10108 3580 1.6 1.0
-personal_finance_tracker original 4306 1482 1.0 1.0
-refactored 7443 2512 1.8 0.6
-in_memory_database original 4565 1671 1.0 1.0
-refactored 6667 2556 1.5 1.0
-command_line_task_manager original 7964 2921 1.0 1.0
-refactored 10387 3899 1.3 1.0
-text_editor original 4324 1323 1.0 1.0
-refactored 5792 1822 1.2 1.0
-concurrent_task_scheduler original 10350 3935 1.0 1.0
-refactored 18653 7165 1.9 0.8
-incremental_backup_system original 4587 1482 1.0 1.0
-refactored 5365 1788 1.3 failed
-Table 6: Full results on MiniCode-repositories large.
-Appendix E Data Generating Prompts
-Prompt to generate general library text descriptions.
-1I need ideas for Python libraries that can be implemented by language models. These libraries should:
-2
-31. Be implementable using only Pythons standard library - no external dependencies
-42. Have enough complexity to demonstrate sophisticated code design (10-20 functions/methods)
-53. Include room for interpretation, so that different implementations can be unique while sharing core functionality
-64. Have clear, practical utility that solves a real programming need
-75. Be realistically implementable by an intelligent language model
-86. Be testable with pytest
-97. Include opportunities for different design approaches (functional vs OOP, etc.)
-10
-11For each library, provide a description that outlines:
-12- The problem domain and core purpose
-13- Key required functionality (without being too prescriptive about implementation details)
-14- Potential use cases that demonstrate practical applications
-15- Suggested extension points where implementers could add their creative spin
-16
-17Please generate several proposals in markdown, following this format:
-18
-19‘‘‘file:<library_name>/DESCRIPTION.md
-20# <Library Name>
-21
-22## Purpose and Motivation
-23<3-5 sentences on what problem this library solves and why its useful>
-24
-25## Core Functionality
-26<Description of 4-6 high-level key features/capabilities without specifying exact implementation>
-27‘‘‘
-28
-29Be creative! Focus on domains where standard Python libraries provide enough building blocks but where a well-designed abstraction layer would add significant value.
-Listing 2: Prompt to generate library descriptions
-1
-2Consider a code repository designed to support the following task description:
-3‘‘‘
-4{task_content}
-5‘‘‘
-6
-7Please list a couple dozen features that would be useful for the repository. Include the specified features as well as several others.
-8
-9List the suggested feature names and descriptions in the following format:
-10
-111: <feature_1_name>: <feature_1_description>
-122: <feature_2_name>: <feature_2_description>
-133: <feature_3_name>: <feature_3_description>
-14...
-1530: <feature_30_name>: <feature_30_description>
-16"""
-17
-18persona_prompt_template = """Consider the following features of a code repository:
-19{listed_features}
-20
-21Think about several possibilities for what kind of person might use this code repository and what they might use it for.
-Please write several brief descriptions for the code repository in first person, formatted in markdown as follows:
-22‘‘‘file:{library_name}/<persona_name>/TASK.md
-23# The Task
-24
-25I am a <...> I want to be able to <...> This code repository <...>
-26
-27# The Requirements
-28
-29* ‘<function_name>‘ : <feature description>
-30* ...
-31‘‘‘
-32
-33Be creative! Write the task description in the style of the proposed persona. Be as exhaustive as possible in including the listed features in the task descriptions requirements.
-Listing 3: Prompt to generate potential uses and features given a library description
-1
-2
-3I need you to implement a Python solution and COMPREHENSIVE suite of tests based on the following task files.
-4Your code must pass the tests provided.
-5
-6{task_content}
-7
-8CRITICAL FORMATTING INSTRUCTIONS:
-91. You MUST format ALL code files exactly as shown below - no exceptions
-102. Start each file with the markdown codeblock marker, followed by "file:" and the relative path
-113. End each file with the closing markdown codeblock marker
-124. Do not use any other format or markdown variations
-135. For test files, do not put them in a subdirectory-- keep them in the outermost level.
-14
-15For each source code file:
-16
-17‘‘‘file:<relative_file_path>
-18<file_content>
-19‘‘‘
-20
-21For each test file:
-22
-23‘‘‘file:<relative_file_path starting with test_>
-24<test_file_content>
-25‘‘‘
-26
-27IMPORTANT:
-28- The opening format must be exactly: ‘‘‘file:path/to/file.py
-29- Do not add language indicators like ‘‘‘python
-30- Do not add explanations between files
-31- Each file must be contained within its own codeblock with the precise format shown above
-32- The system parsing your response requires this exact format to function properly
-33
-34EXAMPLE OUTPUT FORMAT:
-35‘‘‘file:mymodule/mymodule.py
-36def example_function():
-37 return "This is a sample function"
-38‘‘‘
-39
-40‘‘‘file:test_utils.py
-41import pytest
-42from mymodule.mymodule import example_function
-43
-44def test_example_function():
-45 assert example_function() == "This is a sample function"
-46‘‘‘
-47
-48Begin your implementation now, following these formatting rules precisely.
-Listing 4: Prompt to generate initial implementation
-1
-2I need you to fix the implementation of the following code that is failing tests.
-3
-4# Current Implementation:
-5{src_code_content}
-6
-7# Test Files:
-8{test_content}
-9
-10# Failed Tests:
-11{failed_test_details}
-12
-13# Test Output (if available):
-14{test_output}
-15
-16Please carefully analyze the errors and test failures. Pay special attention to:
-171. The exact assertion failures or error messages
-182. What the tests expect vs. what your current implementation provides
-193. Any edge cases or special conditions you might have missed
-20
-21Your task is to fix the implementation to make all tests pass. For each file that needs to be modified,
-22provide the content in the following format:
-23
-24‘‘‘file:<relative_file_path>
-25<file_content>
-26‘‘‘
-27
-28Where <relative_file_path> is the relative path to the file and <file_content> is the updated content of the file.
-29Focus on fixing the specific issues identified in the errors and failed tests while maintaining the overall structure of the code.
-30
-31IMPORTANT: Make targeted changes to address the specific failing test cases. Make sure your implementation passes all test cases,
-32including any edge cases or special conditions mentioned in the tests. Be sure to output code in the specified format.
-Listing 5: Prompt to fix code implementation, given pytest output.
-Appendix F Refactoring examples of Librarian on Code Contests
-F.1 Example 1
+SWE-bench (Jimenez et al., 2024) Low Repository
+Commit-0 (Zhao et al., 2025) Medium Repository
+RefactorBench (Gautam et al., 2025) Low File
+ECCO (Waghjale et al., 2024) High Function
+KernelBench (Ouyang et al., 2025) High Function
+MiniCode(Ours) High Repository
+Appendix I Full MiniCode Codecontests Results
+We present the full agent scores for the CodeContests split in Table 5. The results are given both for each cluster of code sources, as well as averaged across clusters.
+Cluster Agent Tokens $C$ Pass % MDL MDL %
+0 original 9088 95 80.3 11745.85 100.0
+sonnet 3.7 18114 176 87.0 15005.18 127.7
+sonnet 4 11121 138 80.3 9901.53 84.3
+codex-mini 9321 95 80.3 9990.74 85.1
+1 original 12531 255 89.7 13431.86 100.0
+sonnet 3.7 10470 239 96.7 8933.65 66.5
+sonnet 4 11325 298 96.7 8214.42 61.2
+codex-mini 12762 255 89.7 11798.73 87.8
+2 original 14087 376 89.0 15012.77 100.0
+sonnet 3.7 17345 429 91.3 13145.02 87.6
+sonnet 4 14270 356 93.0 10522.66 70.1
+codex-mini 14318 376 89.0 13273.81 88.4
+3 original 14261 246 90.3 13348.82 100.0
+sonnet 3.7 20749 241 97.7 15859.02 118.8
+sonnet 4 13433 197 80.7 11937.04 89.4
+codex-mini 14495 246 90.3 11616.41 87.0
+4 original 17693 336 80.7 14665.16 100.0
+sonnet 3.7 29860 358 100.0 20666.52 141.0
+sonnet 4 18684 352 82.0 12801.21 87.3
+codex-mini 17923 336 80.7 12902.09 88.0
+5 original 12588 286 92.0 12790.11 100.0
+sonnet 3.7 10580 128 99.3 8435.12 65.9
+sonnet 4 10416 155 99.3 9167.85 71.7
+codex-mini 12819 286 92.0 11086.19 86.7
+6 original 11020 131 54.3 13540.41 100.0
+sonnet 3.7 21747 502 88.0 19446.07 143.6
+sonnet 4 11177 143 57.3 10492.00 77.5
+codex-mini 11251 131 54.3 11651.65 86.1
+7 original 12301 180 80.0 12393.73 100.0
+sonnet 3.7 16390 166 91.0 13371.59 107.9
+sonnet 4 11625 150 85.7 9304.25 75.1
+codex-mini 12534 180 80.0 10549.04 85.1
+Avg original 12946 238 82.0 13366.09 100.0
+sonnet 3.7 18157 280 93.9 14357.77 107.4
+sonnet 4 12756 224 84.4 10292.62 77.1
+codex-mini 13178 238 82.0 11608.58 86.8
+Table 5: Comparison of the pass rate and compression metrics of the original files, Claude Sonnet 4 and codex-mini refactorings across CodeContests clusters.
+Appendix J Refactoring examples of Librarian on Code Contests
+J.1 Example 1
 In code snippets LABEL:code:v1, LABEL:helper:v1, LABEL:code:v2, LABEL:helper:v2 one example of 2 refactoring versions.
@@
 71 return res
-Listing 6: Version 1, New Helpers
+Listing 2: Version 1, New Helpers
 1
@@
 52 main()
-Listing 7: Version 1, Refactored Programs
+Listing 3: Version 1, Refactored Programs
 1# ==== NEW HELPER FUNCTIONS ====
@@
 27 return ans
-Listing 8: Version 2, New Helpers
+Listing 4: Version 2, New Helpers
 1# ########## PROGRAM: node_16:cc_python_16 ##########
@@
 66 main()
-Listing 9: Version 2, Refactored Programs
-F.2 Example 2
+Listing 5: Version 2, Refactored Programs
+J.2 Example 2
 In code snippets LABEL:code:v1_2, LABEL:helper:v1_2, LABEL:code:v2_2, LABEL:helper:v2_2 is another example of 2 refactorings where V1 was better according to Librarian.
@@
 6def build_adj_undirected(n, edges):
-7 adj = [[] for _ in range(n)]
+7 adj = $[$] for _ in range(n)]
 8 for u, v, w in edges:
@@
 56 return dist
-Listing 10: Version 1, New Helpers
+Listing 6: Version 1, New Helpers
 1# ########## PROGRAM: node_16:cc_python_16 ##########
@@
 63 # build reversed graph: for each move i->j, add edge j->i
-64 neighbors = [[] for _ in range(n)]
+64 neighbors = $[$] for _ in range(n)]
 65 for i, val in enumerate(a):
@@
 79 main()
-Listing 11: Version 1, Refactored Programs
+Listing 7: Version 1, Refactored Programs
 1#
@@
 67 n = len(a)
-68 go = [[] for _ in range(n)]
+68 go = $[$] for _ in range(n)]
 69 ans = [-1]*n
@@
 86 return ans
-Listing 12: Version 2, New Helpers
+Listing 8: Version 2, New Helpers
 1#
@@
 48 main()
-Listing 13: Version 2, Refactored Programs
-Appendix G Code Agent Failure Modes
-1
-2from datapipe.core import (
-3 tumbling_window,
-4 sliding_window as _sliding_window,
-5 add_serializer,
-6 throttle_upstream as _throttle_upstream,
-7 watermark_event_time as _watermark_event_time,
-8 ...
-9)
-10...
-11
-12def throttle_upstream(max_size):
-13 """
-14 Apply backpressure to slow data ingestion if downstream stages are overloaded.
-15
-16 Args:
-17 max_size: maximum queue size or rate limit
-18
-19 Returns:
-20 Decorator function
-21 """
-22 def decorator(func):
-23 from functools import wraps
-24
-25 @wraps(func)
-26 def queue_wrapper(q, *args, **kwargs):
-27 try:
-28 size = q.qsize()
-29 if size > max_size:
-30 import time
-31 time.sleep(0.01)
-32 except Exception:
-33 pass
-34 return func(q, *args, **kwargs)
-35
-36 return queue_wrapper
-37
-38 return decorator
-39
-40def watermark_event_time(events, allowed_lateness):
-41 """
-42 Assign event-time watermarks to handle late data correctly.
-43
-44 Args:
-45 events: list of dicts with timestamp
-46 allowed_lateness: seconds of allowed lateness
-47
-48 Returns:
-49 Events with watermark annotations
-50 """
-51 # Ensure we return the appropriate format with is_late field
-52 result = []
-53 for e in events:
-54 tagged = dict(e)
-55 max_ts = max(ev[’timestamp’] for ev in events)
-56 watermark = max_ts - allowed_lateness
-57 tagged[’watermark’] = watermark
-58 tagged[’is_late’] = e[’timestamp’] < watermark
-59 result.append(tagged)
-60 return result
-61...
-Listing 14: Claude fails to use imports and instead re-implements the function.
+Listing 9: Version 2, Refactored Programs
+Appendix K Obfuscation example
+We provide the full source code, both refactored and obfuscated, for the MDL and token comparison here.
+Listing 10: Refactored code from modeling_llama.py from the refactored Transformers repository.
+from typing import Optional, Union
+import torch
+from torch import $n$
+from ...cache_utils import Cache, DynamicCache
+from ...generation import GenerationMixin
+from ...masking_utils import create_causal_mask
+from ...modeling_layers import (
+GenericForQuestionAnswering,
+GenericForSequenceClassification,
+GenericForTokenClassification,
+from ...modeling_outputs import (
+BaseModelOutputWithPast,
+CausalLMOutputWithPast,
+from ...modeling_utils import PreTrainedModel
+from ...processing_utils import Unpack
+from ...utils import TransformersKwargs, auto_docstring, can_return_tuple, logging
+from ...utils.generic import check_model_inputs
+from .configuration_llama import LlamaConfig
+from ..shared_library import (
+rotate_half,
+apply_rotary_pos_emb,
+repeat_kv,
+eager_attention_forward,
+RMSNorm,
+BaseMLP,
+BaseRotaryEmbedding,
+BaseAttention,
+BaseDecoderLayer,
+logger = logging.get_logger(__name__)
+class LlamaRMSNorm(RMSNorm):
+pass
+class LlamaRotaryEmbedding(BaseRotaryEmbedding):
+pass
+class LlamaMLP(BaseMLP):
+def __init__(self, config):
+super().__init__(config, mlp_bias=config.mlp_bias)
+class LlamaAttention(BaseAttention):
+def __init__(self, config: LlamaConfig, layer_idx: int):
+super().__init__(
+config=config,
+layer_idx=layer_idx,
+attention_bias=config.attention_bias,
+sliding_window=None
+class LlamaDecoderLayer(BaseDecoderLayer):
+def __init__(self, config: LlamaConfig, layer_idx: int):
+super().__init__(
+config=config,
+layer_idx=layer_idx,
+norm_class=LlamaRMSNorm,
+mlp_class=LlamaMLP,
+attention_class=LlamaAttention
+@auto_docstring
+class LlamaPreTrainedModel(PreTrainedModel):
+config: LlamaConfig
+base_model_prefix = "model"
+supports_gradient_checkpointing = True
+_no_split_modules = ["LlamaDecoderLayer"]
+_skip_keys_device_placement = ["past_key_values"]
+_supports_flash_attn = True
+_supports_sdpa = True
+_supports_flex_attn = True
+_can_compile_fullgraph = True
+_supports_attention_backend = True
+_can_record_outputs = {
+"hidden_states": LlamaDecoderLayer,
+"attentions": LlamaAttention,
+@auto_docstring
+class LlamaModel(LlamaPreTrainedModel):
+def __init__(self, config: LlamaConfig):
+super().__init__(config)
+self.padding_idx = config.pad_token_id
+self.vocab_size = config.vocab_size
+self.embed_tokens = $n$.Embedding(config.vocab_size, config.hidden_size, self.padding_idx)
+self.layers = $n$.ModuleList(
+[LlamaDecoderLayer(config, layer_idx) for layer_idx in range(config.num_hidden_layers)]
+self.norm = LlamaRMSNorm(config.hidden_size, eps=config.rms_norm_eps)
+self.rotary_emb = LlamaRotaryEmbedding(config=config)
+self.gradient_checkpointing = False
+self.post_init()
+@check_model_inputs
+@auto_docstring
+def forward(
+self,
+input_ids: Optional[torch.LongTensor] = None,
+attention_mask: Optional[torch.Tensor] = None,
+position_ids: Optional[torch.LongTensor] = None,
+past_key_values: Optional[Cache] = None,
+inputs_embeds: Optional[torch.FloatTensor] = None,
+cache_position: Optional[torch.LongTensor] = None,
+use_cache: Optional[bool] = None,
+**kwargs: Unpack[TransformersKwargs],
+) -> BaseModelOutputWithPast:
+if (input_ids is None) ^ (inputs_embeds is not None):
+raise ValueError("You must specify exactly one of input_ids or inputs_embeds")
+if inputs_embeds is None:
+inputs_embeds: torch.Tensor = self.embed_tokens(input_ids)
+if use_cache and past_key_values is None:
+past_key_values = DynamicCache(config=self.config)
+if cache_position is None:
+past_seen_tokens = past_key_values.get_seq_length() if past_key_values is not None else 0
+cache_position: torch.Tensor = torch.arange(
+past_seen_tokens, past_seen_tokens + inputs_embeds.shape[1], device=inputs_embeds.device
+if position_ids is None:
+position_ids = cache_position.unsqueeze(0)
+causal_mask = create_causal_mask(
+config=self.config,
+input_embeds=inputs_embeds,
+attention_mask=attention_mask,
+cache_position=cache_position,
+past_key_values=past_key_values,
+position_ids=position_ids,
+hidden_states = inputs_embeds
+position_embeddings = self.rotary_emb(hidden_states, position_ids)
+for decoder_layer in self.layers[: self.config.num_hidden_layers]:
+hidden_states = decoder_layer(
+hidden_states,
+attention_mask=causal_mask,
+position_ids=position_ids,
+past_key_values=past_key_values,
+cache_position=cache_position,
+position_embeddings=position_embeddings,
+**kwargs,
+hidden_states = self.norm(hidden_states)
+return BaseModelOutputWithPast(
+last_hidden_state=hidden_states,
+past_key_values=past_key_values,
+@auto_docstring
+class LlamaForCausalLM(LlamaPreTrainedModel, GenerationMixin):
+_tied_weights_keys = ["lm_head.weight"]
+_tp_plan = {"lm_head": "colwise_rep"}
+_pp_plan = {"lm_head": (["hidden_states"], ["logits"])}
+def __init__(self, config):
+super().__init__(config)
+self.model = LlamaModel(config)
+self.vocab_size = config.vocab_size
+self.lm_head = $n$.Linear(config.hidden_size, config.vocab_size, bias=False)
+self.post_init()
+def set_decoder(self, decoder):
+self.model = decoder
+def get_decoder(self):
+return self.model
+@can_return_tuple
+@auto_docstring
+def forward(
+self,
+input_ids: Optional[torch.LongTensor] = None,
+attention_mask: Optional[torch.Tensor] = None,
+position_ids: Optional[torch.LongTensor] = None,
+past_key_values: Optional[Cache] = None,
+inputs_embeds: Optional[torch.FloatTensor] = None,
+labels: Optional[torch.LongTensor] = None,
+use_cache: Optional[bool] = None,
+cache_position: Optional[torch.LongTensor] = None,
+logits_to_keep: Union[int, torch.Tensor] = 0,
+**kwargs: Unpack[TransformersKwargs],
+) -> CausalLMOutputWithPast:
+r"""
+Example:
+‘‘‘python
+>>> from transformers import AutoTokenizer, LlamaForCausalLM
+>>> model = LlamaForCausalLM.from_pretrained("meta-llama/Llama-2-7b-hf")
+>>> tokenizer = AutoTokenizer.from_pretrained("meta-llama/Llama-2-7b-hf")
+>>> prompt = "Hey, are you conscious? Can you talk to me?"
+>>> inputs = tokenizer(prompt, return_tensors="pt")
+>>> # Generate
+>>> generate_ids = model.generate(inputs.input_ids, max_length=30)
+>>> tokenizer.batch_decode(generate_ids, skip_special_tokens=True, clean_up_tokenization_spaces=False)[0]
+[…]
+outputs: BaseModelOutputWithPast = self.model(
+input_ids=input_ids,
+attention_mask=attention_mask,
+position_ids=position_ids,
+past_key_values=past_key_values,
+inputs_embeds=inputs_embeds,
+use_cache=use_cache,
+cache_position=cache_position,
+**kwargs,
+hidden_states = outputs.last_hidden_state
+slice_indices = slice(-logits_to_keep, None) if isinstance(logits_to_keep, int) else logits_to_keep
+logits = self.lm_head(hidden_states[:, slice_indices, :])
+loss = None
+if labels is not None:
+loss = self.loss_function(logits=logits, labels=labels, vocab_size=self.config.vocab_size, **kwargs)
+return CausalLMOutputWithPast(
+loss=loss,
+logits=logits,
+past_key_values=outputs.past_key_values,
+hidden_states=outputs.hidden_states,
+attentions=outputs.attentions,
+class LlamaForSequenceClassification(GenericForSequenceClassification, LlamaPreTrainedModel): ...
+class LlamaForQuestionAnswering(GenericForQuestionAnswering, LlamaPreTrainedModel):
+base_model_prefix = "transformer"
+class LlamaForTokenClassification(GenericForTokenClassification, LlamaPreTrainedModel): ...
+__all__ = [
+"LlamaForCausalLM",
+"LlamaModel",
+"LlamaPreTrainedModel",
+"LlamaForSequenceClassification",
+"LlamaForQuestionAnswering",
+"LlamaForTokenClassification",
+Listing 11: Obfuscated code from modeling_llama.py from the refactored Transformers repository.
+from typing import Optional as O,Union as U
+import torch as T
+from torch import $n$ as N
+from ...cache_utils import Cache as C,DynamicCache as D
+from ...generation import GenerationMixin as G
+from ...masking_utils import create_causal_mask as m
+from ...modeling_layers import GenericForQuestionAnswering as Q,GenericForSequenceClassification as S,GenericForTokenClassification as K
+from ...modeling_outputs import BaseModelOutputWithPast as B,CausalLMOutputWithPast as L
+from ...modeling_utils import PreTrainedModel as P
+from ...processing_utils import Unpack as X
+from ...utils import TransformersKwargs as W,auto_docstring as a,can_return_tuple as r,logging
+from ...utils.generic import check_model_inputs as c
+from .configuration_llama import LlamaConfig as F
+from ..shared_library import rotate_half,apply_rotary_pos_emb,repeat_kv,eager_attention_forward,RMSNorm as R,BaseMLP as M,BaseRotaryEmbedding as E,BaseAttention as A,BaseDecoderLayer as Y
+l=logging.get_logger(__name__)
+class Z(R):pass
+class I(E):pass
+class J(M):
+def __init__(s,g):super().__init__(g,mlp_bias=g.mlp_bias)
+class H(A):
+def __init__(s,g:F,i:int):super().__init__(config=g,layer_idx=i,attention_bias=g.attention_bias,sliding_window=None)
+class V(Y):
+def __init__(s,g:F,i:int):super().__init__(config=g,layer_idx=i,norm_class=Z,mlp_class=J,attention_class=H)
+@a
+class o(P):
+config:F
+base_model_prefix="model"
+supports_gradient_checkpointing=True
+_no_split_modules=["LlamaDecoderLayer"]
+_skip_keys_device_placement=["past_key_values"]
+_supports_flash_attn=True
+_supports_sdpa=True
+_supports_flex_attn=True
+_can_compile_fullgraph=True
+_supports_attention_backend=True
+_can_record_outputs={"hidden_states":V,"attentions":H}
+@a
+class u(o):
+def __init__(s,g:F):
+super().__init__(g)
+s.padding_idx=g.pad_token_id
+s.vocab_size=g.vocab_size
+s.embed_tokens=N.Embedding(g.vocab_size,g.hidden_size,s.padding_idx)
+s.layers=N.ModuleList([V(g,i)for i in range(g.num_hidden_layers)])
+s.norm=Z(g.hidden_size,eps=g.rms_norm_eps)
+s.rotary_emb=I(config=g)
+s.gradient_checkpointing=False
+s.post_init()
+@c
+@a
+def forward(s,input_ids:O[T.LongTensor]=None,attention_mask:O[T.Tensor]=None,position_ids:O[T.LongTensor]=None,past_key_values:O[C]=None,inputs_embeds:O[T.FloatTensor]=None,cache_position:O[T.LongTensor]=None,use_cache:O[bool]=None,**k:X[W])->B:
+if(input_ids is None)^(inputs_embeds is not None):raise ValueError("You must specify exactly one of input_ids or inputs_embeds")
+if inputs_embeds is None:inputs_embeds:T.Tensor=s.embed_tokens(input_ids)
+if use_cache and past_key_values is None:past_key_values=D(config=s.config)
+if cache_position is None:
+p=past_key_values.get_seq_length()if past_key_values is not None else 0
+cache_position:T.Tensor=T.arange(p,p+inputs_embeds.shape[1],device=inputs_embeds.device)
+if position_ids is None:position_ids=cache_position.unsqueeze(0)
+f=m(config=s.config,input_embeds=inputs_embeds,attention_mask=attention_mask,cache_position=cache_position,past_key_values=past_key_values,position_ids=position_ids)
+h=inputs_embeds
+e=s.rotary_emb(h,position_ids)
+for d in s.layers[:s.config.num_hidden_layers]:h=d(h,attention_mask=f,position_ids=position_ids,past_key_values=past_key_values,cache_position=cache_position,position_embeddings=e,**k)
+h=s.norm(h)
+return B(last_hidden_state=h,past_key_values=past_key_values)
+@a
+class t(o,G):
+_tied_weights_keys=["lm_head.weight"]
+_tp_plan={"lm_head":"colwise_rep"}
+_pp_plan={"lm_head":(["hidden_states"],["logits"])}
+def __init__(s,g):
+super().__init__(g)
+s.model=u(g)
+s.vocab_size=g.vocab_size
+s.lm_head=N.Linear(g.hidden_size,g.vocab_size,bias=False)
+s.post_init()
+def set_decoder(s,d):s.model=d
+def get_decoder(s):return s.model
+@r
+@a
+def forward(s,input_ids:O[T.LongTensor]=None,attention_mask:O[T.Tensor]=None,position_ids:O[T.LongTensor]=None,past_key_values:O[C]=None,inputs_embeds:O[T.FloatTensor]=None,labels:O[T.LongTensor]=None,use_cache:O[bool]=None,cache_position:O[T.LongTensor]=None,logits_to_keep:U[int,T.Tensor]=0,**k:X[W])->L:
+o:B=s.model(input_ids=input_ids,attention_mask=attention_mask,position_ids=position_ids,past_key_values=past_key_values,inputs_embeds=inputs_embeds,use_cache=use_cache,cache_position=cache_position,**k)
+h=o.last_hidden_state
+i=slice(-logits_to_keep,None)if isinstance(logits_to_keep,int)else logits_to_keep
+g=s.lm_head(h[:,i,:])
+n=None
+if labels is not None:n=s.loss_function(logits=g,labels=labels,vocab_size=s.config.vocab_size,**k)
+return L(loss=n,logits=g,past_key_values=o.past_key_values,hidden_states=o.hidden_states,attentions=o.attentions)
+class b(S,o):...
+class x(Q,o):base_model_prefix="transformer"
+class y(K,o):...
+__all__=["t","u","o","b","x","y"]